### v2

Abstract
The Contrastive Language-Image Pretraining (CLIP) model has been widely used in various downstream vision tasks. To augment its capacity for these tasks, the few-shot learning paradigm has been widely adopted. However, current paradigms may struggle with fine-grained classification and satellite image recognition due to domain gaps widening. To address this limitation, we propose retrieval-enhanced visual prompt learning (RePrompt), which introduces retrieval mechanisms to cache the knowledge of downstream tasks. RePrompt constructs a retrieval database from either training examples or external data if available, and uses a retrieval mechanism to enhance multiple stages of a simple prompt learning baseline, thus narrowing the domain gap. During inference, our enhanced model can reference similar samples brought by retrieval to make more accurate predictions. A detailed analysis reveals that retrieval helps to improve the distribution of late features, thus, improving generalization for downstream tasks. Reprompt attains state-of-the-art performance on a wide range of vision datasets, including 11 image datasets, 3 video datasets, 1 multi-view dataset, and 4 domain generalization benchmarks.
Keywords:
few-shot learning prompt tuning image classification video understanding multi-view image classification
1 Introduction
Visual concept recognition has achieved remarkable success in closed-set scenarios with large-scale training sets, typically ImageNet dataset. The recognition accuracy can even surpass human ability. However, it may not always be feasible to have a large training set for each specific visual concept in certain downstream tasks. How to learn a robust visual concept recognition system with low-shot data has become a challenging but valuable problem in computer vision. Previous methods MAML ; prototypical ; matching mainly focus on learning a transferable visual representation from a source domain and quickly adapting to few-shot downstream tasks through fine-tuning techniques. However, limited by the categories and closed-set training samples in the source domain, these few-shot learning algorithms are only effective in impractically simple settings, such as distinguishing $5$-way $1$-shot classification.
Recent advancements in visual representation learning have been propelled by the emergence of large-scale vision-language models, such as CLIP CLIP and ALIGN ALIGN . The feasibility of applying these visual language models in addressing more challenging low-shot learning problems has attracted considerable attention within the research community CLIP ; CoOP ; clip_adapter ; video_prompt . In practice, it is cumbersome to fine-tune the entire vision-language model for transfer learning in each downstream task. Furthermore, finetuning the full model can result in catastrophic forgetting catastrophic_forgetting ; example_forgetting , leading to poor performance on new tasks despite the pre-trained model’s initial success. To address these challenges, CoOP CoOP first proposes a few-shot evaluation framework for visual language models. There are full $C$ classes of downstream tasks and each class has 1/2/4/8/16-shot samples. Meanwhile, CoOP designs a learnable text prompt to replace the sub-optimal hand-crafted text prompt templates.
Nevertheless, the frozen image feature representation also leads to sub-optimal performance. Inspired by VPT VPT , many studies UPT ; MaPLe ; CAVPT ; video_prompt ; promptsrc consider adding additional learnable visual tokens into the image encoder to parameter-efficient fine-tune the vision-language models on downstream tasks. Despite significant improvements achieved by these methods for few-shot learning, these parametric models struggle to generalize to extremely low-shot data or atypical instances, often relying on rote memorization. For instance, given only one image per class, CoOP performs even worse than the zero-shot classification results of CLIP.
Retrieval-augmentation approaches retrieval_long_tail ; RETRO_tokens ; RETROPROMPT ; retrieval_pretraining ; Re-ViLM ; tip-adapter are employed to retrieve knowledge corpus and generate additional references, thereby enhancing performance in low-resource scenarios. Large Language Models (LLM) also get advancements with retrieval-augmentation approaches and further develop Retrieval Augmentation Generation (RAG) RAG . However, the application of retrieval augmentation to visual language models remains underexplored. Tip-Adapter  tip-adapter , a key-value cache model, is constructed as a retrieval auxiliary classifier for CLIP. It leverages the similarity between the test and few-shot training samples, calculating affinity matrixes in text feature space and image feature space. However, it is observed that the variances and means of visual modal similarities are smaller compared to those in visual-language modalities udandarao2022sus . This phenomenon is caused by CLIP maximizing the cosine similarity between paired samples in different modalities during its contrastive learning training, without considering the intra-modality similarity. Consequently, the image similarity metrics employed by Tip-Adapter, derived from CLIP, are potentially compromised by domain gaps and an insufficient adaptation to downstream datasets.
In this work, we investigate the feasibility of learning prompts from a cache model, which facilitates the enhancement of query image features to adjust image-image similarities. To simultaneously optimize these aspects, we introduce Retrieval-enhanced visual Prompt learning (RePrompt), which incorporates a retrieval mechanism for associatively learning knowledge representations from downstream tasks. RePrompt establishes a retrieval database that stores common training image representations as retrieval keys. The current image representations retrieve relevant cached knowledge from the database. Unlike Sus-X udandarao2022sus , which expands the retrieval database using external datasets like LAION schuhmann2022laion , our method adheres strictly to the principles of few-shot learning by not utilizing external datasets. Nonetheless, our method can still benefit from incorporating external knowledge.
To integrate the retrieved knowledge into prompt learning, we consider two enhanced strategies based on the nature of the retrieved value. If the retrieved value is the common training image representation, RePrompt dynamically learns retrieval-enhanced visual prompts based on the retrieved values and inserts them into the inputs of multiple layers of the image encoder. If the retrieved value corresponds to a label, a learnable $k$NN classifier is applied to predict classification results. These results are linearly interpolated into the final prediction. To harmonize these strategies, we utilize a non-parametric $k$NN algorithm between the query instance and the database. This allows us to establish a prior distribution and construct a loss function that guides the training process.
Our experiments show that RePrompt outperforms previous methods and achieves state-of-the-art performance under various few-shot settings for 11 image datasets, 3 video datasets, and a multi-view dataset. Moreover, superior performance on 4 domain generalization benchmarks also demonstrates the generalization ability of our RePrompt for unseen domains.
In summary, the main contributions of our work include:
  • We propose the design of visual Prompt learning with retrieval enhancement, called RePrompt. Specifically, we establish a retrieval database using training examples and implement retrieval-enhanced mechanisms throughout the model, including the input, middle layers, output, and training process.
  • We explore the feasibility of introducing a retrieval system to augment the prompt learning for vision language models, which dynamically select relevant references during inference conditioned on the input. This strategy significantly increases the model performance on downstream few-shot classification
  • The proposed RePrompt has strong flexibility and can be easily extended to tasks other than image datasets. For instance, We extend the framework to video understanding and multi-view recognition tasks.
  • The proposed RePrompt achieves state-of-the-art performance over 11 image datasets, 3 video datasets, and 1 multi-view image dataset, under various few-shot settings. It also demonstrates superior performance on 4 domain generalization benchmarks.
  • We further summarize the patterns between the properties of external memory and retrieval enhancement mechanisms, which will contribute to a better understanding of retrieval enhancement methods in the academic community.
2 Related Work
Retrieval-augmented models have gained significant attention in the fields of Natural Language Processing (NLP) and Computer Vision (CV). Numerous studies have augmented large language models with external memory through retrieval-enhancement mechanism dense_retrieval ; retrieval_pretraining ; bert_knn ; RETRO_tokens ; RETROPROMPT ; retrieval_generation . The integration of knowledge into a language model involves retrieving relevant samples from external memory, thereby enabling the model to generate more informed predictions based on these samples. Typically, the external memory comprises a collection of text paragraphs or a structured knowledge base. The BERT-kNN model bert_knn combines the outputs of a trained BERT model with a non-parametric $k$NN model for question answering tasks. REALM retrieval_pretraining acquires external knowledge from Wikipedia. Karpukhin et al. dense_retrieval developed a dense passage retrieval system for open-domain question answering. RETRO RETRO_tokens demonstrates how to blend retrieved tokens with input tokens in a transformer architecture and methodically evaluates the influence of large-scale external memory datasets on NLP tasks. Chen et al. RETROPROMPT utilize retrieved tokens as input prompts and fine-tune the entire language model. Similar to the approach in RETROPROMPT , we employ retrieval samples as prompts to introduce pertinent information.
Recent advancements in CV also leverage external memory for various tasks. Several semi-parametric methods revisiting_Knn ; generalization_memory ; tip-adapter ; udandarao2022sus ; zhang2023prompt , including Tip-Adapter tip-adapter , have investigated the utilization of a $k$-nearest neighbor (kNN) classifier knn_classifier to enhance classification performance without the need for fine-tuning. RAC retrieval_long_tail integrates the output of a base model, i.e., a conventional vision encoder, with a retrieval module to address long-tail image classification challenges. Blattmann et al. retrieval_diffusion employ a semi-parametric diffusion model for generative image synthesis, which is augmented by a retrieval module.
Building upon the achievements of retrieval augmentation, we introduce a framework for few-shot classification and domain generalization through retrieval-based techniques. Our work focuses on retrieving examples from few-shot training data to construct more dynamic prompts for CLIP models. The most similar method to our own is Training-free Adaption (Tip-Adapter) tip-adapter . Tip-Adapter solely employs retrieval to enhance the final classification distribution. We explore the potential of using retrieval to improve the features extracted by CLIP models.
Finetuning for visual language models is crucial for bridging the domain gap in downstream tasks. While CLIP can perform zero-shot image classification by learning reasonable similarity in high-dimensional joint feature space between image-text pairs, finetuning with few-shot samples in a parameter-efficient manner remains meaningful. To achieve this, CoOP CoOP substitutes the input context token in the text branch of CLIP with learnable parameters prefix for few-shot learning. Following recent advancementsProDA ; ProGrad ; prompt_pretraining , we aim to optimize text prompts more effectively from a gradient perspective. CoCoOP CoCoOP proposes to train an intermediate network to generate image tokens as conditional inputs for the text branch and design a base-to-new benchmark to evaluate the generalization ability of a method within a dataset. Some studies CAVPT ; MaPLe ; promptsrc ; UPT ; video_prompt introduce trainable visual prompts VPT into the vision branch of CLIP to develop the dual prompt tuning scheme. Maple MaPLe proposes to learn layer-wise prompts for the two branches simultaneously. PromptSRC promptsrc utilizes its original features to regularize the learning of prompts.
Adapters are typically subnetworks consisting of two fully connected layers with a nonlinear activation function in between. Referring to the concept of adaptersparameter-efficient , CLIP-Adapter clip_adapter employs lightweight adapters to adapt final features. Tip-Adapter tip-adapter utilizes a key-value cache model for auxiliary classification. Sus-X udandarao2022sus constructs a support set using generative models and extends Tip-Adapter by leveraging image-text affinity matrices. CaFO zhang2023prompt improves visual recognition in low-shot regimes by expanding the training set through a generative model and integrating prior knowledge from multiple pre-training models using multiple adapters.
RePromp provides an in-depth analysis of the key-value cache model from a retrieval perspective. It further explores the potential of this model to contribute to CLIP fine-tuning through retrieval-enhanced visual prompt learning. This approach enables RePrompt to effectively adjust the distribution of the image-feature space utilized in Tip-Adapter.
3 Preliminaries
CLIP CLIP is a pre-trained model that learns aligned vision-language representations from web-scale image-text datasets. It comprises two sub-networks: an image encoder $\mathrm{e}_{\mathrm{I}}$ and a text encoder $\mathrm{e}_{\mathrm{T}}$. The image encoder encodes visual inputs, and the text encoder encodes text inputs, into a shared hidden space $\mathbb{R}^{d}$. Here, $d$ represents the dimension of embeddings (e.g.,d = 512 for ViT ViT in CLIP).
Figure 1: The overall architecture of vision-language prompt tuning. Visual and textual prompt tokens, which are the only learnable parameters in this setup, are incorporated into the vision and language branches of CLIP, respectively. These prompts are designed to optimize performance in low-shot scenarios while preserving the model’s original generalization capabilities.
Classification of CLIP. Given an image $\boldsymbol{x}$ and a set of $C$ category names $T=\{\boldsymbol{t}_{1},\boldsymbol{t}_{2},\dots,\boldsymbol{t}_{C}\}$ (e.g., $C=1000$ for ImageNet imagenet_paper ), the image encoder extracts the image feature $\boldsymbol{z}=e_{\mathrm{I}}(\boldsymbol{x})\in\mathbb{R}^{d}$. The category names in $T$ are combined with a hand-crafted text prompt template “a photo of a [CLASS]” to generate text descriptions. These descriptions are then fed into the text encoder to derive the text features $\boldsymbol{f}\in\mathbb{R}^{d\times C}$. The prediction probability of $\boldsymbol{x}$ belonging to class $c$ is calculated using the inner product similarity between the query image feature $\boldsymbol{z}$ and text features $\boldsymbol{f}$ as shown in Equation 1:
$p(c|x)=\frac{\exp\left(\left(\boldsymbol{z}\cdot\boldsymbol{f}_{c}\right)/\tau\right)}{\sum_{c=1}^{C}\exp\left(\left(\boldsymbol{z}\cdot\boldsymbol{f}_{c}\right)/\tau\right)}.$ (1)
$\tau$ is the temperature coefficient learned by CLIP. The text encoder produces a retrieval-based dynamic classifier.
Text prompt learning. In contrast to hand-crafted prompt engineering, text prompt learning aims to generate more adaptive text features by learning a set of prompts. CoOp CoOP learns a set of parameters $\boldsymbol{P}_{\mathrm{T}}\in\mathbb{R}^{d\times M}$ to replace the predefined prompt templates, where $M$ represents the prompt length. The word token of each category name in $T$ is inserted into a template and treated as the initial values of learnable input prompts. These prompts are further fine-tuned on few-shot data to generate text features. The fine-tuned prompts $\boldsymbol{P}_{\mathrm{T}}$ adjust the decision boundaries of text features for downstream tasks. All parameters inherited from the pretrained CLIP model remain fixed during the training process.
Visual prompt learning. VPT VPT proposes a visual prompt tuning approach, which introduces a few learnable parameters into the input space while freezing the image encoder $e_{\mathrm{I}}$. This method aims to extract more transferable visual features from downstream data. In the context of an image encoder with $L$ layers, the output of the $i$-th layer, $l_{i}$, for $i=1,2,\dots,L$, can be expressed as:
$\left[\boldsymbol{c},\boldsymbol{z}^{i+1}_{1},\dots,\boldsymbol{z}^{i+1}_{S}\right]=l_{i}\left(\boldsymbol{c},\boldsymbol{z}^{i}_{1},\dots,\boldsymbol{z}^{i}_{S}\right),$ (2)
where $\boldsymbol{c}\in\mathbb{R}^{d}$ denotes the classification token and $Z^{i}=\{\boldsymbol{z}^{i}_{1},\dots,\boldsymbol{z}^{i}_{S}\}\in\mathbb{R}^{d\times S}$ represents a set of input image patch tokens of the $i$-th layer, with token sequence length $S$. Furthermore, the learnable visual prompt $P_{\mathrm{I}}\in\mathbb{R}^{d\times N}$ is introduced into the input sequence of the $i$-th layer. $N$ is the length of the visual prompt. There are two visual prompt variants, VPT-Shallow and VPT-Deep. In VPT-Shallow, the class token $\boldsymbol{c}$, along with image patch tokens and visual prompts, is taken as the input of the first layer. VPT-Deep inserts independent visual prompts into each layer.
In this section, we propose a straightforward baseline approach for few-shot image classification, named Vision-Language Prompt Tuning (VLPT). Figure 1 illustrates the overall architecture of VLPT. The visual prompt is optimized through visual prompt learning, while the text prompt is refined via text prompt learning. Upon deriving the image and text features for each category, the model with tuned prompts generates predictions based on the classification paradigm of clip.
Figure 2: The overall workflow of RePrompt includes four main steps: (a) Encoding an image input into a query embedding using a frozen image encoder; (b) Encoding each image entry from the training dataset into key and value embedding pairs with the same frozen image encoder, where the value embeddings also include one-hot representations of labels. We retrieve the top-K relevant knowledge items through maximum inner product search and integrate this knowledge to generate visual prompts; (c) Retrieval-enhanced visual prompts are introduced into the $J$ layers of the visual branch inputs, while the other prompts remain consistent with those of the baseline VLPT; (d) The final output is derived from a linear combination of the prompt-tuned CLIP prediction and the retrieval-aided prediction.
4 Proposed Method
This section provides a detailed explanation of the proposed Retrieval-enhanced Prompt Tuning for VLPT. RePrompt enhances prompt tuning by utilizing the retrieval of relevant information from the training dataset. Given an input image $x$, RePrompt retrieves $k$ potentially useful entries from the retrieval database. Each entry is associated with a corresponding representation $z$ and a label (the construction of the retrieval database is described in Section 4.1). The retriever uses embedding similarity to find relevant downstream knowledge in the retrieval database. We then condition on both the retrieved set and the original input $x$ to generate the output through retrieval-enhanced visual prompt and retrieval-based adapter. Figure 2 illustrates the workflow of RePrompt. Specifically, Section 4.1 describes the process of retrieving entries most relevant to the input query. Subsequently, Section 4.2 details the prompt learner that integrates the query and retrieved knowledge into the prompt learning process. In Section 4.3, we discuss the final prediction in the inference process using the $k$NN-based probability from the retrieval-based adapter. Additionally, this section describes how non-parametric KNN is used between the query and the database to obtain probability distributions, while dynamically controlling the strength of the retrieval mechanism.
4.1 Retrieval module
Our Retrieval module consists of two steps: (1) building the database, and (2) retrieval.
Retrieval database. The retrieval database is constructed by extracting features from the few-shot training dataset $\mathcal{D}$. Specifically, the database contains $\left|\mathcal{D}\right|$ key-value pairs $\left(\boldsymbol{k}_{i},\boldsymbol{v}_{i}\right)$. Each key $\boldsymbol{k}_{i}=\boldsymbol{e}_{\mathrm{R}}(\boldsymbol{x}_{i})\in\mathbb{R}^{d}$ represents the frozen training image representation extracted using $\boldsymbol{e}_{\mathrm{R}}$. Each key-value pair comprises two components: the label $\boldsymbol{y}\in\mathbb{N}^{+}$ and the image representation $\boldsymbol{e}_{\mathrm{R}}(\boldsymbol{x}_{i})\in\mathbb{R}^{d}$. The retrieval database acts as a repository of robust tokens that are adaptable to changes in downstream tasks.
Previous studies retrieval_long_tail ; retrieval_pretraining ; RETRO_tokens have demonstrated the effective performance improvements achieved by expanding the retrieval database. Direct use of external data can disrupt the few-shot task setting. To address this issue, we employ stable diffusion techniques rombach2022high to generate additional training data for various categories. This approach allows us to expand the training data without requiring additional human effort for data collection and annotation. A straightforward template “a photo of a [CLASS]” is utilized as the text prompt. Subsequently, we employ CLIP to filter the top-$\hat{K}$ images of the highest quality for each category. The selected images are added to the original retrieval database, expanding it with new samples and resulting in a total of ($K+\hat{K}$) training images across $C$ categories. To ensure a fair comparison, synthesized data are not included in the training datasets used for the main experiments. In Table 5, we present the performance improvement achieved by incorporating synthesized data, demonstrating an average precision of 84.19% for the 16-shot scenario.
Figure 3: Overview of visual prompt learner and REConv. A visual prompt learner comprises the REConv to generate dynamic visual prompts by learning on retrieved results
Figure 4: Examples of the retrieval mechanism using various encoders $e_{\mathrm{R}}$ on ImageNet under the 16-shot setting. ”ViT-B/16 (Image encoder of CLIP)” indicates that $e_{\mathrm{R}}$ utilizes the ViT-B/16 model as trained by CLIP. ”ViT-B/16 (Timm)” denotes that $e_{\mathrm{R}}$ employs the ViT-B/16 model as trained by Timm with full supervision. ”Top $k$” refers to the $k$ nearest neighbors, and ”IPS” represents the inner product similarity. Employing the image encoder of CLIP as $e_{\mathrm{R}}$ enhances the retrieval of visually similar and more accurate images, particularly for challenging samples such as ”velvet fabric”.
Effective and efficient retrieval. As shown in Figure 2, the retrieval database is represented as a matrix $\boldsymbol{D}\in\mathbb{R}^{\left|\mathcal{D}\right|\times d}$, which serves as a fast approximate $k$-NN of examples. When presented with a query image $\boldsymbol{x}_{\mathrm{q}}$, we utilize the function $e_{\mathrm{R}}$ to map it to a corresponding vector $\boldsymbol{z}_{\mathrm{q}}=\boldsymbol{e}_{\mathrm{R}}(\boldsymbol{x}_{\mathrm{q}})$. Using the query vector $\boldsymbol{z}_{\mathrm{q}}$, we retrieve its approximate $k$-nearest neighbors with representations $\boldsymbol{z}_{1},\boldsymbol{z}_{2},\dots,\boldsymbol{z}_{k}$ from the matrix $\boldsymbol{D}$ based on cosine similarity.
To demonstrate the impact of our retrieval mechanism and the influence of different $e_{\mathrm{R}}$ functions, we include visualizations of retrieved images corresponding to query samples in Figure 4. The visualization results demonstrate the retrieval’s capability to locate similar samples, with the image encoder of CLIP proving to be an effective $e_{\mathrm{R}}$ in identifying more accurately matched images. To ensure an efficient retrieval process, we employ FAISS faiss for querying the database.
4.2 Retrieval-enhanced visual prompt
The proposed method aims to enhance visual prompt learning by utilizing a retrieval database for analogy learning. The visual prompts generated from the retrieval results are referred to as retrieval-enhanced visual prompts.
The retrieval module receives a query vector $\boldsymbol{z}_{\mathrm{q}}$ derived from a raw image $\boldsymbol{x}_{\mathrm{q}}$ and performs a lookup operation in the matrix $\boldsymbol{D}$ to obtain the top $k_{\mathrm{re}}$ most similar candidates. The corresponding representations $\boldsymbol{z}_{1},\boldsymbol{z}_{2},\dots,\boldsymbol{z}_{k_{\mathrm{re}}}$ retrieved from the database are integrated into the image encoder to enhance the visual prompts. Additional fusion vectors $\boldsymbol{z}_{\mathrm{f}}\in\mathbb{R}^{d}$ RETROPROMPT are generated by intuitively aggregating $k_{\mathrm{re}}$ neighboring representations based on their similarity, as follows:
$\boldsymbol{z}_{\mathrm{f}}=\sum_{i=1}^{k_{\mathrm{re}}}\alpha_{i}\cdot\boldsymbol{z}_{i},\quad\alpha_{i}=\frac{e^{\boldsymbol{z}_{\mathrm{q}}\cdot\boldsymbol{z}_{i}}}{\sum_{j=1}^{k_{\mathrm{re}}}e^{\boldsymbol{z}_{\mathrm{q}}\cdot\boldsymbol{z}_{j}}}.$ (3)
The query vectors $\boldsymbol{z}_{\mathrm{q}}$, fusion vectors $\boldsymbol{z}_{\mathrm{f}}$, and retrieved vectors $\boldsymbol{z}_{1},\boldsymbol{z}_{2},\dots,\boldsymbol{z}_{k_{\mathrm{re}}}$ are concatenated to form the input $\hat{\boldsymbol{i}}=\left[\boldsymbol{z}_{\mathrm{q}},\boldsymbol{z}_{\mathrm{f}},\boldsymbol{z}_{1},\dots,\boldsymbol{z}_{k_{\mathrm{re}}}\right]\in\mathbb{R}^{d\times(k_{\mathrm{re}}+2)}$ for the visual prompt learner, which then generates the retrieval-enhanced visual prompts $f_{\mathrm{p}}(\hat{\boldsymbol{i}})\in\mathbb{R}^{d\times(k_{\mathrm{re}}+2)}$.
As depicted at the top of Figure 3, we randomly initialize $J$ visual prompt sets, $\boldsymbol{P}^{1}_{\mathrm{I}},\dots,\boldsymbol{P}^{J}_{\mathrm{I}}$. Subsequently, $J$ retrieval-enhanced convolution (REConv) blocks process $J$ inputs, $\hat{\boldsymbol{i}}_{1},\dots,\hat{\boldsymbol{i}}_{J}$ (by replicating $\hat{\boldsymbol{i}}$ $J$ times), to generate dynamic prompts. These prompts are combined separately with the $J$ visual prompt sets and inserted into the input sequences of the first $J$ layers of the visual branch:
$\left[\boldsymbol{c},\boldsymbol{Z}^{\boldsymbol{j}+1}\right]=l_{\boldsymbol{j}}\left(\boldsymbol{c},f_{\mathrm{p}}(\hat{\boldsymbol{i}}_{\boldsymbol{j}})+\boldsymbol{P}_{\mathrm{I}}^{\boldsymbol{j}},\boldsymbol{Z}^{\boldsymbol{j}}\right),\boldsymbol{j}=1,\dots,J.$ (4)
The remaining $12-J$ layers process learnable prompts without enhancement, and learning these prompts reverts to conventional visual prompt tuning. Furthermore, we discuss the choice of insertion depth in subsection 6.2.
To effectively and efficiently fuse the retrieved representations, we propose a retrieval-enhanced convolution block (REConv). As illustrated at the bottom of Figure 3, a REConv block comprises three convolution layers: two $1\times 1$ convolutions, which individually reduce and scale the channel dimensionality, and a $3\times 3$ convolution positioned between the two $1\times 1$ convolutions. Before these convolution layers, we reshape the 1D token sequence structure of visual prompts into a 2D matrix structure. REConv blocks process the input $\hat{\boldsymbol{i}}$ in parallel to generate dynamic prompts, which can be formulated as
$f_{\mathrm{p}}(\hat{\boldsymbol{i}}_{\boldsymbol{j}})=\beta\mathrm{REconv}_{\boldsymbol{j}}(\mathrm{LN}(\hat{\boldsymbol{i}}_{\boldsymbol{j}}))+\hat{\boldsymbol{i}}_{\boldsymbol{j}},\quad\boldsymbol{j}=1,\dots,J.$ (5)
In this context, $\mathrm{LN}$ denotes Layer Normalization, and $\beta$ is a hyper-parameter used for scaling the output.
4.3 Retrieval-based adapter
We regard Tip-Adapter as a retrieval-based adapter from the perspective of a $k$NN classifier framework. The adapter is trained in conjunction with the retrieval-enhanced visual prompt to enhance the generation of adaptive prediction probabilities for downstream tasks. The query vector, $\hat{\boldsymbol{z}}_{\mathrm{q}}$, is derived from a query instance, ${\boldsymbol{x}}_{\mathrm{q}}$, using an image encoder that integrates retrieval-enhanced visual prompts. This vector retrieves the $\left|\mathcal{D}\right|$ nearest neighbors and their corresponding inner product similarities. We aggregate the probability mass for each label, ${\boldsymbol{y}}_{i}$, across all its occurrences among the retrieved targets. Assuming $\boldsymbol{p}_{k\mathrm{NN}}$ represents the probability of the query instance ${\boldsymbol{x}}_{\mathrm{q}}$ being classified as label $\boldsymbol{y}$, we then reformulate $\boldsymbol{p}_{k\mathrm{NN}}\left(\boldsymbol{y}\mid\boldsymbol{x}_{\mathrm{q}}\right)$ using a weighted sum of the $\boldsymbol{p}_{k\mathrm{NN}}$ probabilities as follows:
$\boldsymbol{p}_{k\mathrm{NN}}\left(\boldsymbol{y}\mid\boldsymbol{x}_{\mathrm{q}}\right)\propto\sum_{\left({\boldsymbol{z}}_{i},{\boldsymbol{y}}_{i}\right)\in\boldsymbol{D}}\mathbf{1}_{{\boldsymbol{y}}={\boldsymbol{y}}_{i}}\exp\left(\hat{{\boldsymbol{z}}}_{\mathrm{q}}\cdot{\boldsymbol{z}}_{i}\right).$ (6)
To better integrate the prediction from the $k$NN and the base model, we interpolate $\boldsymbol{p}\left(\boldsymbol{y}\mid\boldsymbol{x}_{\mathrm{q}}\right)$, blending the $k$NN prediction with the CLIP prediction, scaled by a factor $\lambda$:
$\boldsymbol{p}\left(\boldsymbol{y}\mid\boldsymbol{x}_{\mathrm{q}}\right)=\lambda\boldsymbol{p}_{k\mathrm{NN}}\left(\boldsymbol{y}\mid\boldsymbol{x}_{\mathrm{q}}\right)+(1-\lambda)\boldsymbol{p}_{\mathrm{P}}\left(\boldsymbol{y}\mid\boldsymbol{x}_{\mathrm{q}}\right).$ (7)
Relation to Tip-Adapter. Tip-Adapter constructs an adapter using a key-value cache model from the few-shot training set. The cache model stores prior knowledge encoded in CLIP and facilitates image classification through retrieval in the image embedding space. Tip-Adapter is considered as an auxiliary classifier that leverages image-image similarity. While both approaches aim to improve performance using retrieval, our paper highlights two key distinctions:
  1. (i)
    The features extracted by the CLIP image encoder with retrieval-enhanced prompts are more adaptive to downstream tasks compared to the features obtained by the original CLIP image encoder in Tip-Adapter. As a result, prompt-tuned query features can be utilized to compute improved image-image and image-text similarity matrices for classification.
  2. (ii)
    We extract potential visual context from the cache model and integrate it into the model in the form of prompts, rather than considering the cache model solely as an adapter for classification assistance.
4.4 Retrieval-guiding training
The $k$-nearest neighbor ($k$NN) algorithm primarily focuses on approximating the neighborhoods of query instances knn_classifier . It is intuitive to leverage $k$NN classification results as prior knowledge to guide RePrompt’s focus on hard examples during the training process. Hard samples usually refer to atypical samples with low confidence. To compute a local probability distribution, we limit the number of samples within the retrieved neighbor set $\boldsymbol{K}\subseteq\boldsymbol{D}$, where $k_{rc}\neq\left|\boldsymbol{K}\right|$, as follows:
$\boldsymbol{p}_{k\mathrm{NN}}\left(\boldsymbol{y}\mid\boldsymbol{x}_{\mathrm{q}}\right)\propto\sum_{\left({\boldsymbol{z}}_{i},{\boldsymbol{y}}_{i}\right)\in\boldsymbol{K}}\mathbf{1}_{{\boldsymbol{y}}={\boldsymbol{y}}_{i}}\exp\left({\boldsymbol{z}}_{\mathrm{q}}\cdot{\boldsymbol{k}}_{i}\right).$ (8)
The probability $\boldsymbol{p}_{k\mathrm{NN}}$ represents the confidence of classifying the query instance $\boldsymbol{x}_{\mathrm{q}}$ into specific categories. The negative log-likelihood value of $\boldsymbol{p}_{k\mathrm{NN}}$, similar to Focal Loss focal_loss , serves as the adjustment factor $\boldsymbol{p}_{\mathrm{t}}=-\log\left(\boldsymbol{p}_{k\mathrm{NN}}\right)$. This adjustment factor modifies the relative loss of pseudo-correct or pseudo-error samples distinguished by $k$NN, thereby reweighting the cross-entropy loss $\mathcal{L}_{\mathrm{CE}}$. The final loss is formulated as follows:
$\mathcal{L}=\left(1+\gamma\cdot\boldsymbol{p}_{\mathrm{t}}\right)\mathcal{L}_{\mathrm{CE}},$ (9)
where $\gamma$ is a scaling factor. We set $\left|\boldsymbol{K}\right|=C\times\boldsymbol{n},\boldsymbol{n}\in\mathbb{N}^{+}$ RETROPROMPT , which adopts a similar loss for augmenting the model performance on the few-shot learning task of NLP. In few-shot experiments, $\boldsymbol{n}$ can take values of $1,2,4,8,16$ to accommodate different few-shot settings. We discuss the choice of $n$ in Table 10.
5 Experiments
5.1 Few-shot classification
Datasets. The RePrompt model is evaluated on 11 image classification datasets, covering various tasks such as object recognition (ImageNet1k imagenet_paper and Caltech101  Caltech101_paper ), fine-grained object recognition (Oxford Pets oxford_pest_paper , Flowers102 oxford_flowers_paper , FGVC Aircraft fgvc_aircraft_paper , Stanford Cars Stanford_cars_paper , and Food101 food101_paper ), sense recognition (SUN397 sun397_paper ), remote sensing recognition (EuroSAT eurosat_paper ), texture recognition (DTD DTD_paper ), and action recognition (UCF101 ucf101_paper ). we construct 1,2,4 ,8,16-shot training sets with full categories and the entire test set.
Baseline. We compare our proposed RePrompt with these existing prompt-based approaches: CoOP CoOP , VPT-Deep VPT , Maple MaPLe , PromptSRC promptsrc , Tip-Adapter tip-adapter , Tip-Adapter-F tip-adapter , and VLPT. (1) CoOP learns the context prompt concatenated with $\left[\mathrm{CLASS}\right]$ as the input of text encoder. (2) VPT-Deep incorporates learnable visual prompts into each transformer layer of the visual encoder. (3) VLPT jointly optimizes prompts across different modality encoders, following the methodologies of CoOP and VPT-Deep. (4) Tip-Adapter constructs a cache-model-based adapter using few-shot training data. It is a model which performs few-shot predictions by using those retrieved samples in the end, simply by making some prototypes. (5) Tip-Adapter-F is the variant where the adapter is fine-tuned. (6) Maple learns prompts for the text and image branches simultaneously, rather than a separate side. (7) PromptSRC uses the clip’s original features to regularize the prompt learning.
Training details. Prompt tuning, originally introduced in NLP, is extended to unified visual and language models by implementing it on a ViT model, specifically ViT-B/16, with 12 transformer layers similar to the text encoder. The hyper-parameters of the retrieval modules are set as follows: $k_{\mathrm{re}}$ is set to 7, and 9 random prompts are initialized. The first 7 layers of the vision transformer are equipped with retrieval-enhanced prompts. The prompt and adapter are fine-tuned using the AdamW optimizer with a learning rate of 1e-3 and an epsilon value of 1e-4. We adhere to the data preprocessing protocol of CoOP CoOP and freeze the parameters inherited from the pre-trained model during training.
[…]
Results. The performance of baseline approaches and our proposed RePrompt for few-shot image classification is presented in Figure 5. In general, parameter-efficient finetuning methods perform better than zero-shot classification (zero-shot CLIP) in scenarios with an average of over 11 datasets. RePrompt overall provides consistent improvements on most shots in comparison with all existing methods. RePrompt demonstrates significant performance improvement, particularly with $+5.87\%$ accuracy on DTD and $+9.04\%$ accuracy on Stanford Cars, based on VLPT. RePrompt substantially enhances performance on challenging datasets that contain a wide range of categories, such as ImageNet with 1000 classes and SUN397 with 397 classes. We also observe that Reprompt achieves fewer improvements on Oxford Pets and Food101. This could be attributed to the presence of noisy data and unique data distribution patterns in these datasets CoOP ; food101_paper ; UPT .
5.2 Cross-dataset evaluation
We evaluate the cross-dataset generalization ability of RePrompt by learning on the 16-shot ImageNet setting and then transferring it directly to the remaining 10 datasets. Since this evaluation does not provide training images for other datasets, we remove the retrieval-based adapter module from this test. Table 3 shows the performance comparison between CoOP, CoCoOP, Maple, and PrompSRC. On the ImageNet source dataset, RePrompt achieves the best performance and demonstrates a strong generalization performance by surpassing recent SOTA methods in 5 datasets.
5.3 Domain generalization
Pre-trained vision-language models exhibit a robust ability for domain generalization ability CoOP . We assess the robustness of the proposed RePrompt model on out-of-distribution (OOD) datasets.
Datasets. We follow CoOp CoOP and employ five datasets, namely, ImageNet, ImageNet V2 imagenetv2_paper , ImageNet-Sketch imagenet-sketch_paper , ImageNet-A imagenet-a_paper and ImageNet-R imagenet-r_paper ), to evaluate the generalization ability of RePrompt for out-of-distribution (OOD) data. Following the protocol, we train the model on the ImageNet dataset (as the source dataset) using a 16-shot setting and evaluate its performance on other domain-shifted datasets (as the target datasets). Consequently, we utilize the retrieval database from the ImageNet 16-shot experiment as the retrieval database for the target datasets.
Results. Table 1 summarizes the OOD experiment results, including the accuracy on both source datasets and target datasets. RePrompt achieves optimal results on ImageNet V2 and ImageNet-Sketch while demonstrating performance comparable to PromptSRC promptsrc on ImageNet-R. These results indicate that Reprompt has better generalization for datasets with domain shifts.
Methods ImagNet V2imagenetv2_paper imagenet-sketch_paper imagenet-a_paper Rimagenet-r_paper
CoOP CoOP 71.51 64.20 47.99 49.71 75.21
CoCoOP CoCoOP 71.02 64.07 48.75 50.63 76.18
UPT UPT 72.63 64.35 48.66 50.66 76.24
Maple MaPLe 70.72 64.07 49.15 50.90 76.98
PromptSRC promptsrc 71.27 64.07 49.55 50.90 77.80
RePrompt 74.53 66.66 49.56 49.77 77.48
[…]
Model MVImgNet
$1$-shot $2$-shot $4$-shot $8$-shot $16$-shot Ave.
Vanilla CLIP CLIP 51.4 51.4 51.4 51.4 51.4 51.4
CoOP*  CoOP 62.5 67.8 75.3 79.8 84.0 73.9
VLPT 65.2 69.7 76.8 80.3 84.9 75.3
RePrompt 65.9 73.6 81.1 85.7 90.8 79.4
Table 2: Main results on MVImgNet under the few-shot multi-view settings.RePrompt achieved significant performance improvements due to its retrieval-enhanced module.
Method Source Target Avg.
ImageNet Caltech Oxford Standford Flowers Food101 FGVC SUN397 DTD Euro UCF101
101 Pets Cars 102 Aircraft SAT
CoOPCoOP 71.51 93.70 89.14 64.51 68.71 85.30 18.47 64.15 41.92 46.39 66.55 63.88
CoCoOPCoCoOP 71.02 94.43 90.14 65.32 71.88 86.06 22.94 67.36 45.73 45.37 68.21 65.74
MapleMaPLe 70.72 93.53 90.49 65.57 72.23 86.20 24.74 67.01 46.49 48.06 68.69 66.30
PromptSRCpromptsrc 71.27 93.60 90.25 65.70 70.25 86.15 23.90 67.10 46.87 45.50 68.75 65.81
RePrompt 74.53 94.65 90.16 67.76 72.47 86.00 24.33 67.71 44.39 41.74 69.28 65.85
Table 3: Comparison of RePrompt with existing advanced approaches on cross-dataset benchmark evaluation. The source model (ViT-B/16) is trained on ImageNet imagenet_paper . RePrompt achieves the best results on 5 of 10 datasets. The underline represents the second best.
Model HMDB-51 UCF-101 SSV2
$2$-shot $4$-shot $8$-shot $16$-shot Ave. $2$-shot $4$-shot $8$-shot $16$-shot Ave. $2$-shot $4$-shot $8$-shot $16$-shot Ave.
Vanilla CLIP CLIP 41.9 41.9 41.9 41.9 41.9 63.6 63.6 63.6 63.6 63.6 2.7 2.7 2.7 2.7 2.7
ActionCLIP wang2021actionclip 54.3 56.2 59.3 66.1 59.0 76.7 80.4 87.6 91.8 84.1 4.8 6.9 9.1 12.3 8.3
XCLIP ni2022expanding 60.5 66.8 69.3 71.7 67.1 89.0 91.4 94.7 96.3 92.9 6.6 7.8 9.9 13.7 9.5
A5 ju2022prompting 46.7 50.4 61.3 65.8 56.1 76.3 84.4 90.7 93.0 86.1 4.5 6.7 7.2 9.5 7.0
ViFi-CLIP* rasheed2023fine 61.3 65.3 68.1 70.1 66.4 90.4 92.9 94.4 95.7 93.3 7.2 8.1 10.2 13.9 9.7
RePrompt 63.4 67.6 69.2 71.1 67.8 91.2 93.5 95.0 95.4 93.7 7.6 9.3 9.4 12.1 9.6
Table 4: We compare RePrompt with approaches that explicitly adapt CLIP. ViFi-CLIP only adds an average pooling layer on top of the final features to obtain video-level embeddings. The underline represents the second best.
5.4 Other few-shot classification
We investigate the ability of RePrompt to bridge the modality gap in video and multi-view domains.
Datasets. MVImgNet MVImgNet is a 3D generic dataset that comprises multi-view images, which is a soft bridge between 2D and 3D. It contains 6.5 million frames from 219188 videos, accompanied by comprehensive annotations. These frames covers real-life objects across from 238 classes. Nevertheless, MVImgNet displays a long-tailed distribution of data. Consequently, we eliminate categories with fewer than 16 samples and samples with fewer than 28 frames. In the end, we obtained a subset of MVImgNet, consisting of a total of 220 classes. We adopt the same few-shot setting as used in few-shot image classification. The video understanding task covers HMDB-51 kuehne2011hmdb , UCF-101 ucf101_paper and SSv2 goyal2017ssv2 , where each dataset consists of 2/4/8/16-shot training sets with full categories. The model are evaluated on the first validation split for HMDB-51, UCF-101, and the whole validation split for SSV2.
Training details. We adopt a sparse frame-sampling strategywang2016tsn and configure the number of frames as 16 for multi-view classification and 32 for video understanding. For the video understanding task, models are pre-trained on Kinetics-400k400 to bridge the modality gap and follow a single-view inference rasheed2023fine . he retrieval setup aligns with that of few-shot image classification. However, the training hyperparameters remain consistent with those of ViFi-CLIP-Prompting rasheed2023fine . The minimum retrieval unit in the database is a frame-level embedding, aligned with the input data. The weights of the “Rb adapter” are sourced from another database, where the minimum retrieval unit is a video-level embedding. The use of pooling for simple frame-level temporal aggregation enables the exchange of temporal and multi-view cues in CLIP representations.
Results. Table 2 demonstrates the RePrompt consistently improves performance as the number of shots increases in multi-view classification. For example, it achieves a gain of $+4.2\%$ compared to VLPT. This improvement can be attributed to the retrieval process, which brings more reference information from different perspectives. Table 4 illustrates the effectiveness of RePrompt compared to other CLIP-based approaches for videos. Notably, RePrompt achieves larger gains in low-shot data scenarios, where models are more susceptible to overfitting.
Figure 6: Visualization of attention response map between retrieval-enhanced visual prompts and image patch tokens. The mean self-attention map is from the last vision transformer layers.
5.5 Ablation
Method
EuroSAT
Caltech101
Flowers102
Food101
FGVCAircraft
DTD
OxfordPets
StanfordCars
UCF101
SUN397
ImageNet
average
CoOP* CoOP 80.87 95.50 92.20 86.30 32.57 67.17 92.87 75.03 78.07 73.00 71.20 76.80
VPT-deep*VPT 85.83 95.63 95.00 86.30 41.53 68.93 93.20 76.57 82.90 73.57 70.40 79.80
PromptSRC* 92.30 96.03 97.63 87.47 48.03 73.27 93.73 82.43 86.33 77.27 72.93 82.49
Maple*MaPLe 86.20 95.20 93.57 87.47 37.47 67.53 93.77 74.33 81.20 74.67 71.97 78.49
Tip-Adaptertip-adapter 78.09 95.09 94.62 86.52 39.68 65.90 91.98 75.50 77.93 72.09 70.75 77.10
Tip-Adapter-F*tip-adapter 88.74 95.94 96.79 87.28 46.32 72.87 92.89 84.04 84.38 76.53 73.32 81.74
+VLPT 89.50 95.57 95.37 87.17 48.04 67.83 93.60 76.00 82.90 74.20 71.42 80.14
+Rg training loss 92.48 96.55 94.72 87.56 47.94 69.74 93.54 74.18 83.53 75.20 72.00 80.68
+Re visual prompt 92.60 96.55 96.79 87.57 48.27 70.63 93.62 77.85 83.95 75.60 72.20 81.42
+Rb adapter 92.91 96.51 97.16 87.42 50.32 73.70 93.76 85.04 86.47 77.54 74.53 83.21
CaFO-F*zhang2023prompt 91.72 97.28 97.97 87.15 54.82 71.34 93.89 83.85 86.10 76.62 74.48 83.20
Reprompt+synt. data 91.95 96.80 98.29 87.23 55.21 76.18 93.92 87.55 87.34 77.39 74.23 84.19
Table 5: Component ablation studies over 11 datasets with 16-shot setting. The average accuracy of RePrompt has been steadily improved through the gradual introduction of retrieval enhancement modules.
Figure 7: Visualization of different learned visual feature manifolds via t-SNE. From left to right, we have CLIP, VLPT, and RePrompt. The image features extracted by RePrompt are more discriminating along the z-axis.
Component ablation. We investigate the effectiveness of RePrompt and report results in Table 5. A progressive approach of incorporating retrieval modules is employed to demonstrate the efficacy of enhancement measures at various stages. The average accuracy of RePrompt consistently improves as we incrementally introduce retrieval augmentation: Retrieval-guiding training denoted as $+$Rg training loss, Retrieval-enhanced visual prompt denoted as $+$Re visual prompt, and Retrieval-based adapter denoted as $+$Rb adapter. We also evaluated the Tip-adapter, which solely utilizes the retrieved samples as prototypes. The findings indicate that retrieval-enhanced visual prompts significantly enhance retrieval-based classification results.
[…]
The noticeable improvement achieved with the ”+Rb adapter” can be ascribed to its capability to introduce additional degrees of freedom for the query features, which is unattainable with the conventional Tip-Adapter-F. Tip-Adapter operates as a training-free approach, leveraging the similarity of image features to aid in classification. Despite attempts to mitigate this limitation through fine-tuning in Tip-Adapter-F, the query features remain unchanged. The limited ability of Tip-Adapter-F to manipulate query features significantly hampers its performance.
To further substantiate these observations, we employ t-SNE tSNE to visualize the 3D manifold of test features extracted from the CLIP, VLPT, and RePrompt models on the EuroSAT dataset. The t-SNE visualization results, presented in Figure 7, clearly depict that in the high-dimensional classification space, RePrompt achieves a much more distinct separation of image features, with feature distributions more aligned along the z-axis, distinguishing between various categories. Conversely, Tip-Adapter-F, constrained by a frozen visual encoder, exhibits t-SNE visualization results that closely resemble those of CLIP, with considerable overlap in feature distribution. This overlap elucidates why Tip-Adapter performs suboptimally on remote sensing data such as EuroSAT.
[…]
Discussions on ReConv and retrieved samples. We explore the impact of the network structure, which is responsible for integrating tokens and queries, on the performance outcomes of the ReConv model. To this end, we conduct a supplementary experiment on ImageNet, where we substitute the convolution layer in ReConv with different layers. ReRNN is a variant in which the convolution layer is replaced with an LSTM layer LSTM ; RNN . ReMLP replaces the convolution layer with a multi-layer perceptron layer. Additionally, the quality of high-retrieval samples also contributes to performance; hence, we examine the effect of utilizing original features and random sample features as inputs. ReConv($z_{q}$) is a variant that exclusively uses $z_{q}$ as input, whereas ReConv+Random employs random sample features. The experimental setup for these variants aligns with that of ReConv, maintaining a comparable number of parameters across both models.
As demonstrated in Table 6, ReConv marginally outperforms the other variants across all shot levels. These results suggest that the specific network structure used for token and query fusion may not significantly affect the overall model performance. Instead, the choice of network structure should take into account the computational cost and optimization challenges, particularly as task complexity increases. However, the quality of retrieval samples plays a crucial role, implying that expanding the search library and enhancing the search methodologies are of greater importance.
method 16-shot 8-shot 4-shot 2-shot 1-shot
ReMLP 74.21 73.02 71.40 70.33 69.50
ReRNN 74.27 72.60 71.12 70.32 69.49
ReConv($z_{q}$) 74.34 72.51 70.84 70.42 69.33
ReConv+Random 74.03 72.43 70.68 69.98 66.96
ReConv 74.53 73.36 71.84 70.68 69.87
Table 6: Ablation study of the visual prompt learner across different input strategies and network structures. The first two rows showcase results using RNN and MLP structures, respectively. The middle two rows illustrate the impact of using original image features and random sample features as inputs. ReConv consistently outperforms the other variations, demonstrating its effectiveness across varying shot scenarios.
[…]
Overall, these results highlight the trade-offs between training duration, model complexity, computational efficiency, and performance, emphasizing the effectiveness of RePrompt in leveraging retrieval for improved few-shot classification on high-complexity tasks.
[…]
CoOP CoOP 71.20 0.4 14h40min 200
Tip-Adapter-F tip-adapter 73.32 8.19 5min 20
VLPT 71.42 0.45 15h50min 100
Wo Retrieval 71.91 11.11 3h30min 20
RePrompt 74.53 11.11 4h 20
Table 7: Comparison of training efficiency for different methods on 16-shot ImageNet. All experiments are trained with batch 16 on one RTX309 GPU.
[…]
CoOP CoOP 71.20 299.64 162.5
Tip-Adapter-F tip-adapter 73.32 10.5 42.5
VLPT 71.42 214.84 162.5
Wo Retrieval 71.91 17.57 76.5
RePrompt 74.53 57.60 76.5
Table 8: Comparison of inference efficiency for different methods on 16-shot ImageNet. All experiments are trained with batch 16 and tested with batch size 32 on one RTX309 GPU.
6 Retrieval Discussions
Given that the retrieval database is constructed from training data in a few-shot setting, the performance enhancements observed with RePrompt are intimately linked to data statistics. To delve deeper, we investigate the correlation between the efficacy of retrieval enhancement and the intra-class variance of visual features. This specific relationship is depicted in Figure 8, illustrating the association between intra-class visual variance and performance improvement. Aside from the outliers—“StanfordCars” and “DTD”—the remaining datasets exhibit a discernible pattern. Notably, the labels for StanfordCars correspond to vehicle codes, and those for DTD to texture descriptions, categories where CLIP’s classification capabilities manifest notable deficiencies. Consequently, results from these two datasets are excluded from the general analysis to delineate retrieval enhancement patterns more accurately.
Figure 8: Visualization of the Correlation Between Performance Improvements and Intra-class Visual Variance. The graph demonstrates a general trend where the performance improvements offered by RePrompt tend to decrease as the intra-class visual variance increases. This suggests a sensitivity of the RePrompt’s efficacy to the homogeneity of visual features within classes.
6.1 Analysis of retrieval
Experiments are conducted across FGVC Aircraft, Oxford Pets and ImageNet-1K to quantitatively assess the influence of retrieval in various few-shot scenarios. FGVC Aircraft has the smallest intra-class visual variance, while Oxford Pets has the largest; ImageNet-1K falls intermediate between these two. By comparing the optimal retrieval parameter adjustments among these datasets, we can analysis the influence of intra-class visual variance on the retrieval mechanism.
The increasing value of the retrieval-enhanced prompt number $k_{\mathrm{re}}+2$ indicates that RePrompt benefits from additional retrieved information. As illustrated in Table 9, the optimal $k_{\mathrm{re}}$ parameter varies among the three datasets. Specifically, we observe that $k_{\mathrm{re}}=14$ consistently underperforms across all few-shot settings for Oxford Pets. This suggests that the retrieved information is detrimental and lacks confidence in contexts with significant intra-class visual variance. In contrast, $k_{\mathrm{re}}=14$ demonstrates improved average performance on FGVC Aircraft, underscoring its utility in scenarios with lesser intra-class variance.
$k_{\mathrm{re}}$+2 Dataset 16-shot 8-shot 4-shot 2-shot 1-shot Ave
2+2 FGVC Aircraft 53.05 45.84 39.39 34.86 31.05 40.838
7+2 54.61 47.94 40.05 35.61 31.23 41.888
14+2 54.91 47.97 40.17 35.79 31.17 42.002
2+2 Oxford Pets 94.58 93.84 93.21 92.97 91.55 93.23
7+2 94.85 93.57 93.62 92.94 91.82 93.36
14+2 94.47 93.65 93.16 93.21 91.41 93.18
2+2 ImageNet 74.50 73.23 71.51 70.53 69.46 71.85
7+2 74.53 73.28 71.38 70.40 69.87 71.89
14+2 74.40 73.11 71.29 70.35 69.64 71.76
Table 9: Ablation study on different visual prompt numbers over Fgvc Aircraft and Oxford Pets with few-shot settings. The retrieval-enhanced prompt is mainly affected by the convolution kernel size of REConv.
The factor $n$ determines the number of retrieval samples, $\left|\boldsymbol{K}\right|=C\times\boldsymbol{n}$, used in retrieval-guiding training loss. Increasing the value of $n$ signifies a more relaxed constraint on retrieval. As shown in Table 10, $\boldsymbol{n}=1$ achieves optimal performance in the 16-shot setting on Oxford Pets, which implies that RePrompt prefers the output of CLIP with retrieval-enhanced prompt. However, in low-shot scenarios(e.g., $\boldsymbol{n}=2/1$ under 2/1-shot setting ), RePrompt may require more additional references due to the scarcity of training data.
$\boldsymbol{n}$ Dataset 16-shot 8-shot 4-shot 2-shot 1-shot
16 FGVC Aircraft 54.40 -- -- -- --
8 54.61 47.94 -- -- --
4 54.79 47.28 40.11 -- --
2 54.04 47.31 39.96 35.81 --
1 54.58 47.31 40.05 35.43 31.23
16 Oxford Pets 94.60 -- -- -- --
8 94.49 93.43 -- -- --
4 94.55 93.57 93.62 -- --
2 94.69 93.51 93.21 92.94 --
1 94.85 93.46 93.57 92.70 91.82
16 ImageNet-1k 74.30 -- -- -- --
8 74.53 73.11 -- -- --
4 74.26 73.30 71.52 -- --
2 74.31 73.16 71.50 70.40 --
1 74.37 73.23 71.24 70.38 70.02
Table 10: Ablation studies on retrieving $n$ samples for each class over Fgvc Aircraft and Oxford Pets with few-shot settings. The $k$NN-guiding training is weak in low-shot settings, since the model may require reference on the $k$-nn classifier.
Figure 9: Tuning retrieval parameter results over 3 datasets under the 16-shot learning setting. From top to bottom, images report ablation experiments about tuning retrieval parameters on FGVC Aircraft, ImageNet, and Oxford Pets, respectively.
Based on the above analysis, we summarize two empirical rules as follows: 1) For a dataset with low or high intra-class visual variance, it is imperative to employ a retrieval model with correspondingly strong or weak capabilities to guarantee stable performance improvements across scenarios. 2) In low-shot settings, the inference outcomes are dominantly influenced by the retrieval results.
6.2 Tuning retrieval parameters
The final prediction distribution in RePrompt is influenced by the hyperparameters $\gamma$ and $\lambda$, and the prompt depth $J$. Ablation experiments are conducted to investigate the impact of these hyperparameters on the final outcomes across different datasets.
As illustrated in Figure 9, from top to bottom, ablation experiments are carried out on FGVC Aircraft, ImageNet, and Oxford Pets under the 16-shot learning setting. The experiments adhere to the default configurations, except for modifying the depth $J$, $\gamma$, and $\lambda$ parameters. The experiment data are averaged over the results of experiments on seeds 1,2,3. The reference configuration is marked by a gray vertical dotted line.
[…]
$\gamma$ varies. The hyperparameters $\gamma$ in the guiding loss $\mathcal{L}$ controls the intensity of the retrieval enhancement. As evidenced by the right part of Figure 9, the result continues to improve as $\gamma$ decreases. These results demonstrate that lower values of $\gamma$ enhance the reliance on retrieval-enhanced features, leading to significant improvements in classification accuracy. However, for FGVC Aircraft, the highest accuracy is observed at $\gamma=10^{-1}$, which is caused by its smallest intra-class visual variance.
In summary, the hyper-parameters associated with retrieval critically influence the efficacy of retrieval enhancement. These findings highlight the necessity of striking an optimal balance between incorporating valuable retrieved information and mitigating the potential introduction of noise during the learning process. We provide valuable insights that are instrumental for the design of related methods and establish a guiding framework for future research in this domain.
7 Conclusion
In this paper, we investigate retrieval-enhanced visual prompt learning for visual-language models. Specifically, we propose integrating the results from a retrieval cache model into model inference using two approaches: prompts and predicted distributions. Additionally, We introduce a prompt learner that dynamically adapts prompts based on retrieval results and an auxiliary classification adapter. To regulate the degree of retrieval enhancement, we introduce the prior distribution obtained from semi-parametric retrieval into the cross-entropy loss to guide prompt tuning.
Extensive experiment results prove that the proposed method achieves superior performance over other prompt-learning methods in few-shot learning and comparable results on domain generalization. More significantly, we extend the RePrompt approach from traditional image datasets to more challenging tasks, including video and multi-view image datasets. Finally, we summarize the relationships between the properties of external memory and retrieval enhancement mechanisms. This conclusion stems from a quantitative analysis of the relationship between the hyperparameters of retrieval enhancement and the data distribution of external memory.
We hope that our findings inspire further research in these promising directions: 1) Extending prompt learning to additional downstream tasks, particularly dense prediction tasks such as semantic segmentation and object detection. 2) Investigating the application of retrieval techniques to address challenges, such as long-tail data. Our study lays the groundwork for future research in enhancing visual-language models through retrieval mechanisms, and we encourage further investigations into these critical areas.
8 Data Availability Statements
The datasets that support the findings of this study are openly available in
1) Few-shot image classification: ImageNet, Caltech101, OxfordPets, StanfordCars, Flowers102, Food101, FGVCAircraft, SUN397, DTD, EuroSAT, UCF101. Download split files from this link
2) Domain generalization task: ImageNetV2, ImageNet-Sketch, ImageNet-A, ImageNet-R.
3) Few-shot video understanding: HMDB51, UCF101, Something Something v2 (SSv2), Kinetics
4) Few-shot Multi-few image classification: MVImgNet
Acknowledgments
This work was supported by National Key R&D Program of China (No. 2022ZD0118700) and Natural Science Foundation of Zhejiang Province (NO. LY21F030018).

### v3

@@
-
+Retrieval-Enhanced Visual Prompt Learning for Few-shot Classification
 Abstract
 The Contrastive Language-Image Pretraining (CLIP) model has been widely used in various downstream vision tasks.
-To augment its capacity for these tasks, the few-shot learning paradigm has been widely adopted.
-However, current paradigms may struggle with fine-grained classification and satellite image recognition due to domain gaps widening.
-To address this limitation, we propose retrieval-enhanced visual prompt learning (RePrompt), which introduces retrieval mechanisms to cache the knowledge of downstream tasks.
+The few-shot learning paradigm has been widely adopted to augment its capacity for these tasks.
+However, current paradigms may struggle with fine-grained classification, such as satellite image recognition, due to widening domain gaps.
+To address this limitation, we propose retrieval-enhanced visual prompt learning (RePrompt), which introduces retrieval mechanisms to cache and reuse the knowledge of downstream tasks.
 RePrompt constructs a retrieval database from either training examples or external data if available, and uses a retrieval mechanism to enhance multiple stages of a simple prompt learning baseline, thus narrowing the domain gap.
@@
 Reprompt attains state-of-the-art performance on a wide range of vision datasets, including 11 image datasets, 3 video datasets, 1 multi-view dataset, and 4 domain generalization benchmarks.
-Keywords:
-few-shot learning prompt tuning image classification video understanding multi-view image classification
-1 Introduction
+Index Terms:
+Few-shot learning, prompt tuning, image classification, video understanding, multi-view image classification
+I Introduction
 Visual concept recognition has achieved remarkable success in closed-set scenarios with large-scale training sets, typically ImageNet dataset.
@@
 How to learn a robust visual concept recognition system with low-shot data has become a challenging but valuable problem in computer vision.
-Previous methods MAML ; prototypical ; matching mainly focus on learning a transferable visual representation from a source domain and quickly adapting to few-shot downstream tasks through fine-tuning techniques.
+Previous methods [1, 2, 3] mainly focus on learning a transferable visual representation from a source domain and quickly adapting to few-shot downstream tasks through fine-tuning techniques.
 However, limited by the categories and closed-set training samples in the source domain, these few-shot learning algorithms are only effective in impractically simple settings, such as distinguishing $5$-way $1$-shot classification.
-Recent advancements in visual representation learning have been propelled by the emergence of large-scale vision-language models, such as CLIP CLIP and ALIGN ALIGN .
-The feasibility of applying these visual language models in addressing more challenging low-shot learning problems has attracted considerable attention within the research community CLIP ; CoOP ; clip_adapter ; video_prompt .
+Recent advancements in visual representation learning have been propelled by the emergence of large-scale vision-language models, such as CLIP [4] and ALIGN [5].
+The feasibility of applying these visual language models in addressing more challenging low-shot learning problems has attracted considerable attention within the research community [4, 6, 7, 8].
 In practice, it is cumbersome to fine-tune the entire vision-language model for transfer learning in each downstream task.
-Furthermore, finetuning the full model can result in catastrophic forgetting catastrophic_forgetting ; example_forgetting , leading to poor performance on new tasks despite the pre-trained model’s initial success.
-To address these challenges, CoOP CoOP first proposes a few-shot evaluation framework for visual language models.
+Furthermore, finetuning the full model can result in catastrophic forgetting [9, 10], leading to poor performance on new tasks despite the pre-trained model’s initial success.
+To address these challenges, CoOP [6] first proposes a few-shot evaluation framework for visual language models.
 There are full $C$ classes of downstream tasks and each class has 1/2/4/8/16-shot samples.
@@
 Nevertheless, the frozen image feature representation also leads to sub-optimal performance.
-Inspired by VPT VPT , many studies UPT ; MaPLe ; CAVPT ; video_prompt ; promptsrc consider adding additional learnable visual tokens into the image encoder to parameter-efficient fine-tune the vision-language models on downstream tasks.
+Inspired by VPT [11], many studies [12, 13, 14, 8, 15] consider adding additional learnable visual tokens into the image encoder to parameter-efficient fine-tune the vision-language models on downstream tasks.
 Despite significant improvements achieved by these methods for few-shot learning, these parametric models struggle to generalize to extremely low-shot data or atypical instances, often relying on rote memorization.
 For instance, given only one image per class, CoOP performs even worse than the zero-shot classification results of CLIP.
-Retrieval-augmentation approaches retrieval_long_tail ; RETRO_tokens ; RETROPROMPT ; retrieval_pretraining ; Re-ViLM ; tip-adapter are employed to retrieve knowledge corpus and generate additional references, thereby enhancing performance in low-resource scenarios.
-Large Language Models (LLM) also get advancements with retrieval-augmentation approaches and further develop Retrieval Augmentation Generation (RAG) RAG .
+Retrieval-augmentation approaches [16, 17, 18, 19, 20, 21] are employed to retrieve knowledge corpus and generate additional references, thereby enhancing performance in low-resource scenarios.
+Large Language Models (LLM) also get advancements with retrieval-augmentation approaches and further develop Retrieval Augmentation Generation (RAG) [22].
 However, the application of retrieval augmentation to visual language models remains underexplored.
-Tip-Adapter  tip-adapter , a key-value cache model, is constructed as a retrieval auxiliary classifier for CLIP.
+Tip-Adapter [21], a key-value cache model, is constructed as a retrieval auxiliary classifier for CLIP.
 It leverages the similarity between the test and few-shot training samples, calculating affinity matrixes in text feature space and image feature space.
-However, it is observed that the variances and means of visual modal similarities are smaller compared to those in visual-language modalities udandarao2022sus .
+However, it is observed that the variances and means of visual modal similarities are smaller compared to those in visual-language modalities [23].
 This phenomenon is caused by CLIP maximizing the cosine similarity between paired samples in different modalities during its contrastive learning training, without considering the intra-modality similarity.
@@
 In this work, we investigate the feasibility of learning prompts from a cache model, which facilitates the enhancement of query image features to adjust image-image similarities.
-To simultaneously optimize these aspects, we introduce Retrieval-enhanced visual Prompt learning (RePrompt), which incorporates a retrieval mechanism for associatively learning knowledge representations from downstream tasks.
+To simultaneously optimize these aspects, we introduce Retrieval-enhanced Visual Prompt learning (RePrompt), which incorporates a retrieval mechanism for learning knowledge representations from downstream tasks.
 RePrompt establishes a retrieval database that stores common training image representations as retrieval keys.
 The current image representations retrieve relevant cached knowledge from the database.
-Unlike Sus-X udandarao2022sus , which expands the retrieval database using external datasets like LAION schuhmann2022laion , our method adheres strictly to the principles of few-shot learning by not utilizing external datasets.
+Unlike Sus-X [23], which expands the retrieval database using external datasets like LAION [24], our method adheres strictly to the principles of few-shot learning by not utilizing external datasets.
 Nonetheless, our method can still benefit from incorporating external knowledge.
@@
   • We further summarize the patterns between the properties of external memory and retrieval enhancement mechanisms, which will contribute to a better understanding of retrieval enhancement methods in the academic community.
-2 Related Work
+II Related Work
 Retrieval-augmented models have gained significant attention in the fields of Natural Language Processing (NLP) and Computer Vision (CV).
-Numerous studies have augmented large language models with external memory through retrieval-enhancement mechanism dense_retrieval ; retrieval_pretraining ; bert_knn ; RETRO_tokens ; RETROPROMPT ; retrieval_generation .
+Numerous studies have augmented large language models with external memory through retrieval-enhancement mechanism [25, 19, 26, 17, 18, 27].
 The integration of knowledge into a language model involves retrieving relevant samples from external memory, thereby enabling the model to generate more informed predictions based on these samples.
 Typically, the external memory comprises a collection of text paragraphs or a structured knowledge base.
-The BERT-kNN model bert_knn combines the outputs of a trained BERT model with a non-parametric $k$NN model for question answering tasks.
-REALM retrieval_pretraining acquires external knowledge from Wikipedia.
+The BERT-kNN model [26] combines the outputs of a trained BERT model with a non-parametric $k$NN model for question-answering tasks.
+REALM [19] acquires external knowledge from Wikipedia.
 Karpukhin et al.
-dense_retrieval developed a dense passage retrieval system for open-domain question answering.
-RETRO RETRO_tokens demonstrates how to blend retrieved tokens with input tokens in a transformer architecture and methodically evaluates the influence of large-scale external memory datasets on NLP tasks.
+[25] developed a dense passage retrieval system for open-domain question answering.
+RETRO [17] demonstrates how to blend retrieved tokens with input tokens in a transformer architecture and methodically evaluates the influence of large-scale external memory datasets on NLP tasks.
 Chen et al.
-RETROPROMPT utilize retrieved tokens as input prompts and fine-tune the entire language model.
-Similar to the approach in RETROPROMPT , we employ retrieval samples as prompts to introduce pertinent information.
+[18] utilize retrieved tokens as input prompts and fine-tune the entire language model.
+Similar to the approach in [18], we employ retrieval samples as prompts to introduce pertinent information.
 Recent advancements in CV also leverage external memory for various tasks.
-Several semi-parametric methods revisiting_Knn ; generalization_memory ; tip-adapter ; udandarao2022sus ; zhang2023prompt , including Tip-Adapter tip-adapter , have investigated the utilization of a $k$-nearest neighbor (kNN) classifier knn_classifier to enhance classification performance without the need for fine-tuning.
-RAC retrieval_long_tail integrates the output of a base model, i.e., a conventional vision encoder, with a retrieval module to address long-tail image classification challenges.
+Several semi-parametric methods [28, 29, 21, 23, 30], including Tip-Adapter [21], have investigated the utilization of a $k$-nearest neighbor (kNN) classifier [31] to enhance classification performance without the need for fine-tuning.
+RAC [16] integrates the output of a base model, i.e., a conventional vision encoder, with a retrieval module to address long-tail image classification challenges.
 Blattmann et al.
-retrieval_diffusion employ a semi-parametric diffusion model for generative image synthesis, which is augmented by a retrieval module.
+[32] employ a semi-parametric diffusion model for generative image synthesis, which is augmented by a retrieval module.
 Building upon the achievements of retrieval augmentation, we introduce a framework for few-shot classification and domain generalization through retrieval-based techniques.
 Our work focuses on retrieving examples from few-shot training data to construct more dynamic prompts for CLIP models.
-The most similar method to our own is Training-free Adaption (Tip-Adapter) tip-adapter .
+The most similar method to our own is Training-free Adaption (Tip-Adapter) [21].
 Tip-Adapter solely employs retrieval to enhance the final classification distribution.
@@
 While CLIP can perform zero-shot image classification by learning reasonable similarity in high-dimensional joint feature space between image-text pairs, finetuning with few-shot samples in a parameter-efficient manner remains meaningful.
-To achieve this, CoOP CoOP substitutes the input context token in the text branch of CLIP with learnable parameters prefix for few-shot learning.
-Following recent advancementsProDA ; ProGrad ; prompt_pretraining , we aim to optimize text prompts more effectively from a gradient perspective.
-CoCoOP CoCoOP proposes to train an intermediate network to generate image tokens as conditional inputs for the text branch and design a base-to-new benchmark to evaluate the generalization ability of a method within a dataset.
-Some studies CAVPT ; MaPLe ; promptsrc ; UPT ; video_prompt introduce trainable visual prompts VPT into the vision branch of CLIP to develop the dual prompt tuning scheme.
-Maple MaPLe proposes to learn layer-wise prompts for the two branches simultaneously.
-PromptSRC promptsrc utilizes its original features to regularize the learning of prompts.
+To achieve this, CoOP [6] substitutes the input context token in the text branch of CLIP with learnable parameters [33] for few-shot learning.
+Following recent advancements [34, 35, 36], we aim to optimize text prompts more effectively from a gradient perspective.
+CoCoOP [37] proposes to train an intermediate network to generate image tokens as conditional inputs for the text branch and design a base-to-new benchmark to evaluate the generalization ability of a method within a dataset.
+Some studies [14, 13, 15, 12, 8] introduce trainable visual prompts[11] into the vision branch of CLIP to develop the dual prompt tuning scheme.
+Maple [13, 38] proposes to learn layer-wise prompts for the two branches simultaneously.
+PromptSRC [15] utilizes its original features to regularize the learning of prompts.
+Beyond their focus on few-shot classification tasks, learnable prompts exhibit remarkable adaptability across diverse downstream applications, offering a unified and efficient framework for fine-tuning foundation models to address various complex tasks [39, 40, 41, 42, 43].
 Adapters are typically subnetworks consisting of two fully connected layers with a nonlinear activation function in between.
-Referring to the concept of adaptersparameter-efficient , CLIP-Adapter clip_adapter employs lightweight adapters to adapt final features.
-Tip-Adapter tip-adapter utilizes a key-value cache model for auxiliary classification.
-Sus-X udandarao2022sus constructs a support set using generative models and extends Tip-Adapter by leveraging image-text affinity matrices.
-CaFO zhang2023prompt improves visual recognition in low-shot regimes by expanding the training set through a generative model and integrating prior knowledge from multiple pre-training models using multiple adapters.
+Referring to the concept of adapters [44], CLIP-Adapter [7] employs lightweight adapters to adapt final features.
+Tip-Adapter‘[21] utilizes a key-value cache model for auxiliary classification.
+DPT [38] propose class-aware visual prompts to guide the CLIP model focus on the target object in downstream work.
+Sus-X [23] constructs a support set using generative models and extends Tip-Adapter by leveraging image-text affinity matrices.
+CaFO [30] improves visual recognition in low-shot regimes by expanding the training set through a generative model and integrating prior knowledge from multiple pre-training models using multiple adapters.
 RePromp provides an in-depth analysis of the key-value cache model from a retrieval perspective.
@@
 This approach enables RePrompt to effectively adjust the distribution of the image-feature space utilized in Tip-Adapter.
-3 Preliminaries
-CLIP CLIP is a pre-trained model that learns aligned vision-language representations from web-scale image-text datasets.
+III Preliminaries
+CLIP [4] is a pre-trained model that learns aligned vision-language representations from web-scale image-text datasets.
 It comprises two sub-networks: an image encoder $\mathrm{e}_{\mathrm{I}}$ and a text encoder $\mathrm{e}_{\mathrm{T}}$.
 The image encoder encodes visual inputs, and the text encoder encodes text inputs, into a shared hidden space $\mathbb{R}^{d}$.
-Here, $d$ represents the dimension of embeddings (e.g.,d = 512 for ViT ViT in CLIP).
+Here, $d$ represents the dimension of embeddings (e.g.,d = 512 for ViT[45] in CLIP).
 Figure 1: The overall architecture of vision-language prompt tuning.
@@
 Classification of CLIP.
-Given an image $\boldsymbol{x}$ and a set of $C$ category names $T=\{\boldsymbol{t}_{1},\boldsymbol{t}_{2},\dots,\boldsymbol{t}_{C}\}$ (e.g., $C=1000$ for ImageNet imagenet_paper ), the image encoder extracts the image feature $\boldsymbol{z}=e_{\mathrm{I}}(\boldsymbol{x})\in\mathbb{R}^{d}$.
+Given an image $\boldsymbol{x}$ and a set of $C$ category names $T=\{\boldsymbol{t}_{1},\boldsymbol{t}_{2},\dots,\boldsymbol{t}_{C}\}$ (e.g., $C=1000$ for ImageNet [46]), the image encoder extracts the image feature $\boldsymbol{z}=e_{\mathrm{I}}(\boldsymbol{x})\in\mathbb{R}^{d}$.
 The category names in $T$ are combined with a hand-crafted text prompt template “a photo of a [CLASS]” to generate text descriptions.
 These descriptions are then fed into the text encoder to derive the text features $\boldsymbol{f}\in\mathbb{R}^{d\times C}$.
-The prediction probability of $\boldsymbol{x}$ belonging to class $c$ is calculated using the inner product similarity between the query image feature $\boldsymbol{z}$ and text features $\boldsymbol{f}$ as shown in Equation 1:
-$p(c|x)=\frac{\exp\left(\left(\boldsymbol{z}\cdot\boldsymbol{f}_{c}\right)/\tau\right)}{\sum_{c=1}^{C}\exp\left(\left(\boldsymbol{z}\cdot\boldsymbol{f}_{c}\right)/\tau\right)}.$ (1)
-$\tau$ is the temperature coefficient learned by CLIP. The text encoder produces a retrieval-based dynamic classifier.
+The prediction probability of $\boldsymbol{x}$ belonging to class $c$ is calculated using the inner product similarity between the query image feature $\boldsymbol{z}$ and text features $\boldsymbol{f}$ as following:
+$p(c|x)=\frac{\exp\left(\left(\boldsymbol{z}\cdot\boldsymbol{f}_{c}\right)/\tau\right)}{\sum_{c=1}^{C}\exp\left(\left(\boldsymbol{z}\cdot\boldsymbol{f}_{c}\right)/\tau\right)},$ (1)
+where $\tau$ is the temperature coefficient learned by CLIP. The text encoder produces a retrieval-based dynamic classifier.
 Text prompt learning.
 In contrast to hand-crafted prompt engineering, text prompt learning aims to generate more adaptive text features by learning a set of prompts.
-CoOp CoOP learns a set of parameters $\boldsymbol{P}_{\mathrm{T}}\in\mathbb{R}^{d\times M}$ to replace the predefined prompt templates, where $M$ represents the prompt length.
+CoOp [6] learns a set of parameters $\boldsymbol{P}_{\mathrm{T}}\in\mathbb{R}^{d\times M}$ to replace the predefined prompt templates, where $M$ represents the prompt length.
 The word token of each category name in $T$ is inserted into a template and treated as the initial values of learnable input prompts.
@@
 Visual prompt learning.
-VPT VPT proposes a visual prompt tuning approach, which introduces a few learnable parameters into the input space while freezing the image encoder $e_{\mathrm{I}}$.
+VPT [11] proposes a visual prompt tuning approach, which introduces a few learnable parameters into the input space while freezing the image encoder $e_{\mathrm{I}}$.
 This method aims to extract more transferable visual features from downstream data.
@@
 In this section, we propose a straightforward baseline approach for few-shot image classification, named Vision-Language Prompt Tuning (VLPT).
-Figure 1 illustrates the overall architecture of VLPT.
+Fig.
+1 illustrates the overall architecture of VLPT.
 The visual prompt is optimized through visual prompt learning, while the text prompt is refined via text prompt learning.
@@
 We retrieve the top-K relevant knowledge items through maximum inner product search and integrate this knowledge to generate visual prompts; (c) Retrieval-enhanced visual prompts are introduced into the $J$ layers of the visual branch inputs, while the other prompts remain consistent with those of the baseline VLPT; (d) The final output is derived from a linear combination of the prompt-tuned CLIP prediction and the retrieval-aided prediction.
-4 Proposed Method
+The retrieval database for each dataset comes from a corresponding few-shot training set for fairness.
+IV Proposed Method
 This section provides a detailed explanation of the proposed Retrieval-enhanced Prompt Tuning for VLPT.
@@
 Given an input image $x$, RePrompt retrieves $k$ potentially useful entries from the retrieval database.
-Each entry is associated with a corresponding representation $z$ and a label (the construction of the retrieval database is described in Section 4.1).
+Each entry is associated with a corresponding representation $z$ and a label (the construction of the retrieval database is described in Section IV-A).
 The retriever uses embedding similarity to find relevant downstream knowledge in the retrieval database.
-We then condition on both the retrieved set and the original input $x$ to generate the output through retrieval-enhanced visual prompt and retrieval-based adapter.
-Figure 2 illustrates the workflow of RePrompt.
-Specifically, Section 4.1 describes the process of retrieving entries most relevant to the input query.
-Subsequently, Section 4.2 details the prompt learner that integrates the query and retrieved knowledge into the prompt learning process.
-In Section 4.3, we discuss the final prediction in the inference process using the $k$NN-based probability from the retrieval-based adapter.
+We then condition both the retrieved set and the original input $x$ to generate the output through the retrieval-enhanced visual prompt and retrieval-based adapter.
+Fig.
+2 illustrates the workflow of RePrompt.
+Specifically, Section IV-A describes the process of retrieving entries most relevant to the input query.
+Subsequently, Section IV-B details the prompt learner that integrates the query and retrieved knowledge into the prompt learning process.
+In Section IV-C, we discuss the final prediction in the inference process using the $k$NN-based probability from the retrieval-based adapter.
 Additionally, this section describes how non-parametric KNN is used between the query and the database to obtain probability distributions, while dynamically controlling the strength of the retrieval mechanism.
-4.1 Retrieval module
-Our Retrieval module consists of two steps: (1) building the database, and (2) retrieval.
-Retrieval database.
+IV-A Retrieval Module
+Our retrieval mechanism/process consists of two steps: (1) retrieval database construction, and (2) nearest neighbor retrieval.
+Retrieval database construction.
 The retrieval database is constructed by extracting features from the few-shot training dataset $\mathcal{D}$.
 Specifically, the database contains $\left|\mathcal{D}\right|$ key-value pairs $\left(\boldsymbol{k}_{i},\boldsymbol{v}_{i}\right)$.
-Each key $\boldsymbol{k}_{i}=\boldsymbol{e}_{\mathrm{R}}(\boldsymbol{x}_{i})\in\mathbb{R}^{d}$ represents the frozen training image representation extracted using $\boldsymbol{e}_{\mathrm{R}}$.
-Each key-value pair comprises two components: the label $\boldsymbol{y}\in\mathbb{N}^{+}$ and the image representation $\boldsymbol{e}_{\mathrm{R}}(\boldsymbol{x}_{i})\in\mathbb{R}^{d}$.
+Each key $\boldsymbol{k}_{i}=\boldsymbol{e}_{\mathrm{R}}(\boldsymbol{x}_{i})\in\mathbb{R}^{d}$ represents the training image representation encoded by the frozen image encoder $\boldsymbol{e}_{\mathrm{R}}$.
+In our method, values of each key-value pair comprise two components: the label $\boldsymbol{y}\in\mathbb{N}^{+}$ (one-hot encoding) and the image representation $\boldsymbol{e}_{\mathrm{R}}(\boldsymbol{x}_{i})\in\mathbb{R}^{d}$ encoded by $\boldsymbol{e}_{\mathrm{R}}$.
 The retrieval database acts as a repository of robust tokens that are adaptable to changes in downstream tasks.
-Previous studies retrieval_long_tail ; retrieval_pretraining ; RETRO_tokens have demonstrated the effective performance improvements achieved by expanding the retrieval database.
+Previous studies [16, 19, 17] have demonstrated the effective performance improvements achieved by expanding the retrieval database.
 Direct use of external data can disrupt the few-shot task setting.
-To address this issue, we employ stable diffusion techniques rombach2022high to generate additional training data for various categories.
-This approach allows us to expand the training data without requiring additional human effort for data collection and annotation.
-A straightforward template “a photo of a [CLASS]” is utilized as the text prompt.
-Subsequently, we employ CLIP to filter the top-$\hat{K}$ images of the highest quality for each category.
-The selected images are added to the original retrieval database, expanding it with new samples and resulting in a total of ($K+\hat{K}$) training images across $C$ categories.
-To ensure a fair comparison, synthesized data are not included in the training datasets used for the main experiments.
-In Table 5, we present the performance improvement achieved by incorporating synthesized data, demonstrating an average precision of 84.19% for the 16-shot scenario.
-Figure 3: Overview of visual prompt learner and REConv. A visual prompt learner comprises the REConv to generate dynamic visual prompts by learning on retrieved results
-Figure 4: Examples of the retrieval mechanism using various encoders $e_{\mathrm{R}}$ on ImageNet under the 16-shot setting.
-”ViT-B/16 (Image encoder of CLIP)” indicates that $e_{\mathrm{R}}$ utilizes the ViT-B/16 model as trained by CLIP.
-”ViT-B/16 (Timm)” denotes that $e_{\mathrm{R}}$ employs the ViT-B/16 model as trained by Timm with full supervision.
-”Top $k$” refers to the $k$ nearest neighbors, and ”IPS” represents the inner product similarity.
-Employing the image encoder of CLIP as $e_{\mathrm{R}}$ enhances the retrieval of visually similar and more accurate images, particularly for challenging samples such as ”velvet fabric”.
-Effective and efficient retrieval.
-As shown in Figure 2, the retrieval database is represented as a matrix $\boldsymbol{D}\in\mathbb{R}^{\left|\mathcal{D}\right|\times d}$, which serves as a fast approximate $k$-NN of examples.
-When presented with a query image $\boldsymbol{x}_{\mathrm{q}}$, we utilize the function $e_{\mathrm{R}}$ to map it to a corresponding vector $\boldsymbol{z}_{\mathrm{q}}=\boldsymbol{e}_{\mathrm{R}}(\boldsymbol{x}_{\mathrm{q}})$.
+To ensure a fair comparison, synthesized data are not included in the retrieval database used for main experiments.
+In the ablation study, we employ a stable diffusion generation model [47] to generate additional training data to expand training data without requiring additional human effort for data collection and annotation.
+The synthesized images are added to the original retrieval database.
+In Table V, we present the performance improvement achieved by incorporating synthesized data, demonstrating an average precision of 84.19% for the 16-shot scenario.
+Nearest neighbor retrieval.
+As shown in Fig.
+2, the retrieval database is represented as a matrix $\boldsymbol{D}\in\mathbb{R}^{\left|\mathcal{D}\right|\times d}$, which serves as a fast approximate $k$-NN of examples.
+Given a query image $\boldsymbol{x}_{\mathrm{q}}$, we utilize the encoder $e_{\mathrm{R}}$ to map it to a corresponding vector $\boldsymbol{z}_{\mathrm{q}}=\boldsymbol{e}_{\mathrm{R}}(\boldsymbol{x}_{\mathrm{q}})$.
 Using the query vector $\boldsymbol{z}_{\mathrm{q}}$, we retrieve its approximate $k$-nearest neighbors with representations $\boldsymbol{z}_{1},\boldsymbol{z}_{2},\dots,\boldsymbol{z}_{k}$ from the matrix $\boldsymbol{D}$ based on cosine similarity.
-To demonstrate the impact of our retrieval mechanism and the influence of different $e_{\mathrm{R}}$ functions, we include visualizations of retrieved images corresponding to query samples in Figure 4.
-The visualization results demonstrate the retrieval’s capability to locate similar samples, with the image encoder of CLIP proving to be an effective $e_{\mathrm{R}}$ in identifying more accurately matched images.
-To ensure an efficient retrieval process, we employ FAISS faiss for querying the database.
-4.2 Retrieval-enhanced visual prompt
+The image encoder of CLIP is proved to be an effective $e_{\mathrm{R}}$ in identifying more accurately matched images.
+To ensure an efficient retrieval process, we employ FAISS [48] for querying the database.
+IV-B Retrieval-enhanced Visual Prompting
 The proposed method aims to enhance visual prompt learning by utilizing a retrieval database for analogy learning.
@@
 The corresponding representations $\boldsymbol{z}_{1},\boldsymbol{z}_{2},\dots,\boldsymbol{z}_{k_{\mathrm{re}}}$ retrieved from the database are integrated into the image encoder to enhance the visual prompts.
-Additional fusion vectors $\boldsymbol{z}_{\mathrm{f}}\in\mathbb{R}^{d}$ RETROPROMPT are generated by intuitively aggregating $k_{\mathrm{re}}$ neighboring representations based on their similarity, as follows:
+Additional fusion vectors $\boldsymbol{z}_{\mathrm{f}}\in\mathbb{R}^{d}$ [18] are generated by intuitively aggregating $k_{\mathrm{re}}$ neighboring representations based on their similarity, as follows:
 $\boldsymbol{z}_{\mathrm{f}}=\sum_{i=1}^{k_{\mathrm{re}}}\alpha_{i}\cdot\boldsymbol{z}_{i},\quad\alpha_{i}=\frac{e^{\boldsymbol{z}_{\mathrm{q}}\cdot\boldsymbol{z}_{i}}}{\sum_{j=1}^{k_{\mathrm{re}}}e^{\boldsymbol{z}_{\mathrm{q}}\cdot\boldsymbol{z}_{j}}}.$ (3)
 The query vectors $\boldsymbol{z}_{\mathrm{q}}$, fusion vectors $\boldsymbol{z}_{\mathrm{f}}$, and retrieved vectors $\boldsymbol{z}_{1},\boldsymbol{z}_{2},\dots,\boldsymbol{z}_{k_{\mathrm{re}}}$ are concatenated to form the input $\hat{\boldsymbol{i}}=\left[\boldsymbol{z}_{\mathrm{q}},\boldsymbol{z}_{\mathrm{f}},\boldsymbol{z}_{1},\dots,\boldsymbol{z}_{k_{\mathrm{re}}}\right]\in\mathbb{R}^{d\times(k_{\mathrm{re}}+2)}$ for the visual prompt learner, which then generates the retrieval-enhanced visual prompts $f_{\mathrm{p}}(\hat{\boldsymbol{i}})\in\mathbb{R}^{d\times(k_{\mathrm{re}}+2)}$.
-As depicted at the top of Figure 3, we randomly initialize $J$ visual prompt sets, $\boldsymbol{P}^{1}_{\mathrm{I}},\dots,\boldsymbol{P}^{J}_{\mathrm{I}}$.
+Figure 3: Overview of visual prompt learner and REConv. A visual prompt learner comprises the REConv to generate dynamic visual prompts by learning on retrieved results
+As depicted at the top of Fig.
+3, we randomly initialize $J$ visual prompt sets, $\boldsymbol{P}^{1}_{\mathrm{I}},\dots,\boldsymbol{P}^{J}_{\mathrm{I}}$.
 Subsequently, $J$ retrieval-enhanced convolution (REConv) blocks process $J$ inputs, $\hat{\boldsymbol{i}}_{1},\dots,\hat{\boldsymbol{i}}_{J}$ (by replicating $\hat{\boldsymbol{i}}$ $J$ times), to generate dynamic prompts.
@@
 $\left[\boldsymbol{c},\boldsymbol{Z}^{\boldsymbol{j}+1}\right]=l_{\boldsymbol{j}}\left(\boldsymbol{c},f_{\mathrm{p}}(\hat{\boldsymbol{i}}_{\boldsymbol{j}})+\boldsymbol{P}_{\mathrm{I}}^{\boldsymbol{j}},\boldsymbol{Z}^{\boldsymbol{j}}\right),\boldsymbol{j}=1,\dots,J.$ (4)
+The output token corresponding to the visual prompt is discarded.
 The remaining $12-J$ layers process learnable prompts without enhancement, and learning these prompts reverts to conventional visual prompt tuning.
-Furthermore, we discuss the choice of insertion depth in subsection 6.2.
-To effectively and efficiently fuse the retrieved representations, we propose a retrieval-enhanced convolution block (REConv).
-As illustrated at the bottom of Figure 3, a REConv block comprises three convolution layers: two $1\times 1$ convolutions, which individually reduce and scale the channel dimensionality, and a $3\times 3$ convolution positioned between the two $1\times 1$ convolutions.
+We propose a retrieval-enhanced convolution block to effectively and efficiently fuse the retrieved representations (REConv).
+As illustrated at the bottom of Fig.
+3, a REConv block comprises three convolution layers: two $1\times 1$ convolutions, which individually reduce and scale the channel dimensionality and a $3\times 3$ convolution positioned between the two $1\times 1$ convolutions.
 Before these convolution layers, we reshape the 1D token sequence structure of visual prompts into a 2D matrix structure.
@@
 In this context, $\mathrm{LN}$ denotes Layer Normalization, and $\beta$ is a hyper-parameter used for scaling the output.
-4.3 Retrieval-based adapter
+IV-C Retrieval-based Adapter
 We regard Tip-Adapter as a retrieval-based adapter from the perspective of a $k$NN classifier framework.
@@
 Tip-Adapter is considered as an auxiliary classifier that leverages image-image similarity.
-While both approaches aim to improve performance using retrieval, our paper highlights two key distinctions:
-  1. (i)
-    The features extracted by the CLIP image encoder with retrieval-enhanced prompts are more adaptive to downstream tasks compared to the features obtained by the original CLIP image encoder in Tip-Adapter.
+While both approaches aim to improve performance using retrieval, our paper highlights two key distinctions: (1) The features extracted by the CLIP image encoder with retrieval-enhanced prompts are more adaptive to downstream tasks compared to the features obtained by the original CLIP image encoder in Tip-Adapter.
 As a result, prompt-tuned query features can be utilized to compute improved image-image and image-text similarity matrices for classification.
-  2. (ii)
-    We extract potential visual context from the cache model and integrate it into the model in the form of prompts, rather than considering the cache model solely as an adapter for classification assistance.
-4.4 Retrieval-guiding training
-The $k$-nearest neighbor ($k$NN) algorithm primarily focuses on approximating the neighborhoods of query instances knn_classifier .
+(2) We extract potential visual context from the cache model and integrate it into the model in the form of prompts, rather than considering the cache model solely as an adapter for classification assistance.
+IV-D Retrieval-guided Training
+The $k$-nearest neighbor ($k$NN) algorithm primarily focuses on approximating the neighborhoods of query instances[31].
 It is intuitive to leverage $k$NN classification results as prior knowledge to guide RePrompt’s focus on hard examples during the training process.
@@
 The probability $\boldsymbol{p}_{k\mathrm{NN}}$ represents the confidence of classifying the query instance $\boldsymbol{x}_{\mathrm{q}}$ into specific categories.
-The negative log-likelihood value of $\boldsymbol{p}_{k\mathrm{NN}}$, similar to Focal Loss focal_loss , serves as the adjustment factor $\boldsymbol{p}_{\mathrm{t}}=-\log\left(\boldsymbol{p}_{k\mathrm{NN}}\right)$.
+The negative log-likelihood value of $\boldsymbol{p}_{k\mathrm{NN}}$, similar to Focal Loss[49], serves as the adjustment factor $\boldsymbol{p}_{\mathrm{t}}=-\log\left(\boldsymbol{p}_{k\mathrm{NN}}\right)$.
 This adjustment factor modifies the relative loss of pseudo-correct or pseudo-error samples distinguished by $k$NN, thereby reweighting the cross-entropy loss $\mathcal{L}_{\mathrm{CE}}$.
@@
 where $\gamma$ is a scaling factor.
-We set $\left|\boldsymbol{K}\right|=C\times\boldsymbol{n},\boldsymbol{n}\in\mathbb{N}^{+}$ RETROPROMPT , which adopts a similar loss for augmenting the model performance on the few-shot learning task of NLP.
-In few-shot experiments, $\boldsymbol{n}$ can take values of $1,2,4,8,16$ to accommodate different few-shot settings.
-We discuss the choice of $n$ in Table 10.
-5 Experiments
-5.1 Few-shot classification
+We set $\left|\boldsymbol{K}\right|=C\times\boldsymbol{n},\boldsymbol{n}\in\mathbb{N}^{+}$[18], which adopts a similar loss for augmenting the model performance on the few-shot learning task of NLP.
+In few-shot experiments, $\boldsymbol{n}$ can take values of $1,2,4,8,16$ to accommodate few-shot settings.
+We discuss the choice of $n$ in Table X.
+V Experiments
+[…]
+V-A Few-shot Classification
 Datasets.
-The RePrompt model is evaluated on 11 image classification datasets, covering various tasks such as object recognition (ImageNet1k imagenet_paper and Caltech101  Caltech101_paper ), fine-grained object recognition (Oxford Pets oxford_pest_paper , Flowers102 oxford_flowers_paper , FGVC Aircraft fgvc_aircraft_paper , Stanford Cars Stanford_cars_paper , and Food101 food101_paper ), sense recognition (SUN397 sun397_paper ), remote sensing recognition (EuroSAT eurosat_paper ), texture recognition (DTD DTD_paper ), and action recognition (UCF101 ucf101_paper ).
+The RePrompt model is evaluated on 11 image classification datasets, covering various tasks such as object recognition (ImageNet1k [46] and Caltech101 [50]), fine-grained object recognition (Oxford Pets [51], Flowers102 [52], FGVC Aircraft [53], Stanford Cars [54], and Food101 [55]), sense recognition (SUN397 [56]), remote sensing recognition (EuroSAT [57]), texture recognition (DTD [58]), and action recognition (UCF101 [59]).
 we construct 1,2,4 ,8,16-shot training sets with full categories and the entire test set.
 Baseline.
-We compare our proposed RePrompt with these existing prompt-based approaches: CoOP CoOP , VPT-Deep VPT , Maple MaPLe , PromptSRC promptsrc , Tip-Adapter tip-adapter , Tip-Adapter-F tip-adapter , and VLPT.
+We compare our proposed RePrompt with these existing prompt-based approaches: CoOP [6], VPT-Deep [11], Maple [13], PromptSRC [15], Tip-Adapter [21], Tip-Adapter-F [21], and VLPT.
 (1) CoOP learns the context prompt concatenated with $\left[\mathrm{CLASS}\right]$ as the input of text encoder.
@@
 (4) Tip-Adapter constructs a cache-model-based adapter using few-shot training data.
-It is a model which performs few-shot predictions by using those retrieved samples in the end, simply by making some prototypes.
+It is a model that performs few-shot predictions by using those retrieved samples in the end, simply by making some prototypes.
 (5) Tip-Adapter-F is the variant where the adapter is fine-tuned.
@@
 The prompt and adapter are fine-tuned using the AdamW optimizer with a learning rate of 1e-3 and an epsilon value of 1e-4.
-We adhere to the data preprocessing protocol of CoOP CoOP and freeze the parameters inherited from the pre-trained model during training.
-[…]
+We adhere to the data preprocessing protocol of CoOP and freeze the parameters inherited from the pre-trained model during training.
 Results.
-The performance of baseline approaches and our proposed RePrompt for few-shot image classification is presented in Figure 5.
+The performance of baseline approaches and our proposed RePrompt for few-shot image classification is presented in Fig.
+4.
 In general, parameter-efficient finetuning methods perform better than zero-shot classification (zero-shot CLIP) in scenarios with an average of over 11 datasets.
@@
 We also observe that Reprompt achieves fewer improvements on Oxford Pets and Food101.
-This could be attributed to the presence of noisy data and unique data distribution patterns in these datasets CoOP ; food101_paper ; UPT .
-5.2 Cross-dataset evaluation
+This could be attributed to the noisy data and unique data distribution patterns in these datasets [6, 55, 12].
+Cross-dataset Evaluation.
 We evaluate the cross-dataset generalization ability of RePrompt by learning on the 16-shot ImageNet setting and then transferring it directly to the remaining 10 datasets.
 Since this evaluation does not provide training images for other datasets, we remove the retrieval-based adapter module from this test.
-Table 3 shows the performance comparison between CoOP, CoCoOP, Maple, and PrompSRC.
+Table I shows the performance comparison between CoOP, CoCoOP, Maple, and PrompSRC.
 On the ImageNet source dataset, RePrompt achieves the best performance and demonstrates a strong generalization performance by surpassing recent SOTA methods in 5 datasets.
-5.3 Domain generalization
-Pre-trained vision-language models exhibit a robust ability for domain generalization ability CoOP . We assess the robustness of the proposed RePrompt model on out-of-distribution (OOD) datasets.
+The lower performance on FGVC Aircraft, DTD, and UCF101 can be attributed to the different nature of their tasks, which require more specialized feature extraction.
+However, we adopt the same retrieval database sampled from ImageNet.
+V-B Domain Generalization
+Pre-trained vision-language models exhibit a robust ability for domain generalization ability [6]. We assess the robustness of the proposed RePrompt model on out-of-distribution (OOD) datasets.
 Datasets.
-We follow CoOp CoOP and employ five datasets, namely, ImageNet, ImageNet V2 imagenetv2_paper , ImageNet-Sketch imagenet-sketch_paper , ImageNet-A imagenet-a_paper and ImageNet-R imagenet-r_paper ), to evaluate the generalization ability of RePrompt for out-of-distribution (OOD) data.
+We follow CoOp [6] and employ five datasets, namely, ImageNet, ImageNet V2 [60], ImageNet-Sketch [61], ImageNet-A [62] and ImageNet-R [63]), to evaluate the generalization ability of RePrompt for out-of-distribution (OOD) data.
 Following the protocol, we train the model on the ImageNet dataset (as the source dataset) using a 16-shot setting and evaluate its performance on other domain-shifted datasets (as the target datasets).
@@
 Results.
-Table 1 summarizes the OOD experiment results, including the accuracy on both source datasets and target datasets.
-RePrompt achieves optimal results on ImageNet V2 and ImageNet-Sketch while demonstrating performance comparable to PromptSRC promptsrc on ImageNet-R.
+Table II summarizes the OOD experiment results, including the accuracy on both source datasets and target datasets.
+RePrompt achieves optimal results on ImageNet V2 and ImageNet-Sketch while demonstrating performance comparable to PromptSRC [15] on ImageNet-R.
 These results indicate that Reprompt has better generalization for datasets with domain shifts.
-Methods ImagNet V2imagenetv2_paper imagenet-sketch_paper imagenet-a_paper Rimagenet-r_paper
-CoOP CoOP 71.51 64.20 47.99 49.71 75.21
-CoCoOP CoCoOP 71.02 64.07 48.75 50.63 76.18
-UPT UPT 72.63 64.35 48.66 50.66 76.24
-Maple MaPLe 70.72 64.07 49.15 50.90 76.98
-PromptSRC promptsrc 71.27 64.07 49.55 50.90 77.80
-RePrompt 74.53 66.66 49.56 49.77 77.48
-[…]
-Model MVImgNet
-$1$-shot $2$-shot $4$-shot $8$-shot $16$-shot Ave.
-Vanilla CLIP CLIP 51.4 51.4 51.4 51.4 51.4 51.4
-CoOP*  CoOP 62.5 67.8 75.3 79.8 84.0 73.9
-VLPT 65.2 69.7 76.8 80.3 84.9 75.3
-RePrompt 65.9 73.6 81.1 85.7 90.8 79.4
-Table 2: Main results on MVImgNet under the few-shot multi-view settings.RePrompt achieved significant performance improvements due to its retrieval-enhanced module.
+RePrompt struggles on ImageNet-A and ImageNet-R.
+This is likely due to a lack of diversity in the retrieval database, only ImageNet, which limits generalization to adversarial and artistic images.
+In practical scenarios, we can improve the deficiency by expanding the search base and improving the search quality.
+TABLE I: Comparison of RePrompt with existing advanced approaches on cross-dataset benchmark evaluation.
+The source model (ViT-B/16) is trained on ImageNet[46].
+RePrompt achieves the best results on 5 of 10 datasets.
+The underline represents the second best.
 Method Source Target Avg.
@@
 101 Pets Cars 102 Aircraft SAT
-CoOPCoOP 71.51 93.70 89.14 64.51 68.71 85.30 18.47 64.15 41.92 46.39 66.55 63.88
-CoCoOPCoCoOP 71.02 94.43 90.14 65.32 71.88 86.06 22.94 67.36 45.73 45.37 68.21 65.74
-MapleMaPLe 70.72 93.53 90.49 65.57 72.23 86.20 24.74 67.01 46.49 48.06 68.69 66.30
-PromptSRCpromptsrc 71.27 93.60 90.25 65.70 70.25 86.15 23.90 67.10 46.87 45.50 68.75 65.81
+CoOP[6] 71.51 93.70 89.14 64.51 68.71 85.30 18.47 64.15 41.92 46.39 66.55 63.88
+CoCoOP[37] 71.02 94.43 90.14 65.32 71.88 86.06 22.94 67.36 45.73 45.37 68.21 65.74
+Maple[13] 70.72 93.53 90.49 65.57 72.23 86.20 24.74 67.01 46.49 48.06 68.69 66.30
+PromptSRC[15] 71.27 93.60 90.25 65.70 70.25 86.15 23.90 67.10 46.87 45.50 68.75 65.81
 RePrompt 74.53 94.65 90.16 67.76 72.47 86.00 24.33 67.71 44.39 41.74 69.28 65.85
-Table 3: Comparison of RePrompt with existing advanced approaches on cross-dataset benchmark evaluation.
-The source model (ViT-B/16) is trained on ImageNet imagenet_paper .
-RePrompt achieves the best results on 5 of 10 datasets.
+[…]
+Methods ImagNet V2[60] S[61] A[62] R[63]
+CoOP[6] 71.51 64.20 47.99 49.71 75.21
+CoCoOP[37] 71.02 64.07 48.75 50.63 76.18
+UPT[12] 72.63 64.35 48.66 50.66 76.24
+Maple[13] 70.72 64.07 49.15 50.90 76.98
+PromptSRC[15] 71.27 64.07 49.55 50.90 77.80
+RePrompt 74.53 66.66 49.56 49.77 77.48
+TABLE III: The main results on MVImgNet are under the few-shot multi-view settings.RePrompt achieved significant performance improvements due to its retrieval-enhanced module.
+The methods marked with “*” are reproduced by us.
+Model MVImgNet
+$1$-shot $2$-shot $4$-shot $8$-shot $16$-shot Ave.
+Vanilla CLIP[4] 51.4 51.4 51.4 51.4 51.4 51.4
+CoOP* [6] 62.5 67.8 75.3 79.8 84.0 73.9
+VLPT 65.2 69.7 76.8 80.3 84.9 75.3
+RePrompt 65.9 73.6 81.1 85.7 90.8 79.4
+TABLE IV: We compare RePrompt with approaches that explicitly adapt CLIP.
+ViFi-CLIP only adds an average pooling layer on top of the final features to obtain video-level embeddings.
 The underline represents the second best.
@@
 $2$-shot $4$-shot $8$-shot $16$-shot Ave. $2$-shot $4$-shot $8$-shot $16$-shot Ave. $2$-shot $4$-shot $8$-shot $16$-shot Ave.
-Vanilla CLIP CLIP 41.9 41.9 41.9 41.9 41.9 63.6 63.6 63.6 63.6 63.6 2.7 2.7 2.7 2.7 2.7
-ActionCLIP wang2021actionclip 54.3 56.2 59.3 66.1 59.0 76.7 80.4 87.6 91.8 84.1 4.8 6.9 9.1 12.3 8.3
-XCLIP ni2022expanding 60.5 66.8 69.3 71.7 67.1 89.0 91.4 94.7 96.3 92.9 6.6 7.8 9.9 13.7 9.5
-A5 ju2022prompting 46.7 50.4 61.3 65.8 56.1 76.3 84.4 90.7 93.0 86.1 4.5 6.7 7.2 9.5 7.0
-ViFi-CLIP* rasheed2023fine 61.3 65.3 68.1 70.1 66.4 90.4 92.9 94.4 95.7 93.3 7.2 8.1 10.2 13.9 9.7
+Vanilla CLIP[4] 41.9 41.9 41.9 41.9 41.9 63.6 63.6 63.6 63.6 63.6 2.7 2.7 2.7 2.7 2.7
+ActionCLIP[64] 54.3 56.2 59.3 66.1 59.0 76.7 80.4 87.6 91.8 84.1 4.8 6.9 9.1 12.3 8.3
+XCLIP[65] 60.5 66.8 69.3 71.7 67.1 89.0 91.4 94.7 96.3 92.9 6.6 7.8 9.9 13.7 9.5
+A5[66] 46.7 50.4 61.3 65.8 56.1 76.3 84.4 90.7 93.0 86.1 4.5 6.7 7.2 9.5 7.0
+ViFi-CLIP*[67] 61.3 65.3 68.1 70.1 66.4 90.4 92.9 94.4 95.7 93.3 7.2 8.1 10.2 13.9 9.7
 RePrompt 63.4 67.6 69.2 71.1 67.8 91.2 93.5 95.0 95.4 93.7 7.6 9.3 9.4 12.1 9.6
-Table 4: We compare RePrompt with approaches that explicitly adapt CLIP.
-ViFi-CLIP only adds an average pooling layer on top of the final features to obtain video-level embeddings.
-The underline represents the second best.
-5.4 Other few-shot classification
+V-C Additional Few-shot Classification
 We investigate the ability of RePrompt to bridge the modality gap in video and multi-view domains.
 Datasets.
-MVImgNet MVImgNet is a 3D generic dataset that comprises multi-view images, which is a soft bridge between 2D and 3D.
+MVImgNet [68] is a 3D generic dataset that comprises multi-view images, which is a soft bridge between 2D and 3D.
 It contains 6.5 million frames from 219188 videos, accompanied by comprehensive annotations.
@@
 We adopt the same few-shot setting as used in few-shot image classification.
-The video understanding task covers HMDB-51 kuehne2011hmdb , UCF-101 ucf101_paper and SSv2 goyal2017ssv2 , where each dataset consists of 2/4/8/16-shot training sets with full categories.
+The video understanding task covers HMDB-51 [69], UCF-101 [59] and SSv2 [70], where each dataset consists of 2/4/8/16-shot training sets with full categories.
 The model are evaluated on the first validation split for HMDB-51, UCF-101, and the whole validation split for SSV2.
 Training details.
-We adopt a sparse frame-sampling strategywang2016tsn and configure the number of frames as 16 for multi-view classification and 32 for video understanding.
-For the video understanding task, models are pre-trained on Kinetics-400k400 to bridge the modality gap and follow a single-view inference rasheed2023fine .
+We adopt a sparse frame-sampling strategy [71] and configure the number of frames as 16 for multi-view classification and 32 for video understanding.
+For the video understanding task, models are pre-trained on Kinetics-400 [72] to bridge the modality gap and follow a single-view inference [67].
 he retrieval setup aligns with that of few-shot image classification.
-However, the training hyperparameters remain consistent with those of ViFi-CLIP-Prompting rasheed2023fine .
+However, the training hyperparameters remain consistent with those of ViFi-CLIP-Prompting [67].
 The minimum retrieval unit in the database is a frame-level embedding, aligned with the input data.
@@
 Results.
-Table 2 demonstrates the RePrompt consistently improves performance as the number of shots increases in multi-view classification.
+Table III demonstrates the RePrompt consistently improves performance as the number of shots increases in multi-view classification.
 For example, it achieves a gain of $+4.2\%$ compared to VLPT.
 This improvement can be attributed to the retrieval process, which brings more reference information from different perspectives.
-Table 4 illustrates the effectiveness of RePrompt compared to other CLIP-based approaches for videos.
+Table IV illustrates the effectiveness of RePrompt compared to other CLIP-based approaches for videos.
 Notably, RePrompt achieves larger gains in low-shot data scenarios, where models are more susceptible to overfitting.
 Figure 5: Visualization of attention response map between retrieval-enhanced visual prompts and image patch tokens. The mean self-attention map is from the last vision transformer layers.
-5.5 Ablation
+V-D Ablation Study
+TABLE V: Component ablation studies over 11 datasets with 16-shot setting.
+The average accuracy of RePrompt has been steadily improved through the gradual introduction of retrieval enhancement modules.
+We reproduce methods marked with “*” on current datasets.
+“Reprompt+Synthetic data” means introducing synthetic data on the basis of RePrompt.
 Method
@@
 average
-CoOP* CoOP 80.87 95.50 92.20 86.30 32.57 67.17 92.87 75.03 78.07 73.00 71.20 76.80
-VPT-deep*VPT 85.83 95.63 95.00 86.30 41.53 68.93 93.20 76.57 82.90 73.57 70.40 79.80
+CoOP*[6] 80.87 95.50 92.20 86.30 32.57 67.17 92.87 75.03 78.07 73.00 71.20 76.80
+VPT-deep*[11] 85.83 95.63 95.00 86.30 41.53 68.93 93.20 76.57 82.90 73.57 70.40 79.80
 PromptSRC* 92.30 96.03 97.63 87.47 48.03 73.27 93.73 82.43 86.33 77.27 72.93 82.49
-Maple*MaPLe 86.20 95.20 93.57 87.47 37.47 67.53 93.77 74.33 81.20 74.67 71.97 78.49
-Tip-Adaptertip-adapter 78.09 95.09 94.62 86.52 39.68 65.90 91.98 75.50 77.93 72.09 70.75 77.10
-Tip-Adapter-F*tip-adapter 88.74 95.94 96.79 87.28 46.32 72.87 92.89 84.04 84.38 76.53 73.32 81.74
+Maple*[13] 86.20 95.20 93.57 87.47 37.47 67.53 93.77 74.33 81.20 74.67 71.97 78.49
+Tip-Adapter[21] 78.09 95.09 94.62 86.52 39.68 65.90 91.98 75.50 77.93 72.09 70.75 77.10
+Tip-Adapter-F*[21] 88.74 95.94 96.79 87.28 46.32 72.87 92.89 84.04 84.38 76.53 73.32 81.74
 +VLPT 89.50 95.57 95.37 87.17 48.04 67.83 93.60 76.00 82.90 74.20 71.42 80.14
@@
 +Re visual prompt 92.60 96.55 96.79 87.57 48.27 70.63 93.62 77.85 83.95 75.60 72.20 81.42
-+Rb adapter 92.91 96.51 97.16 87.42 50.32 73.70 93.76 85.04 86.47 77.54 74.53 83.21
-CaFO-F*zhang2023prompt 91.72 97.28 97.97 87.15 54.82 71.34 93.89 83.85 86.10 76.62 74.48 83.20
-Reprompt+synt. data 91.95 96.80 98.29 87.23 55.21 76.18 93.92 87.55 87.34 77.39 74.23 84.19
-Table 5: Component ablation studies over 11 datasets with 16-shot setting.
-The average accuracy of RePrompt has been steadily improved through the gradual introduction of retrieval enhancement modules.
++Rb adapter(RePrompt) 92.91 96.51 97.16 87.42 50.32 73.70 93.76 85.04 86.47 77.54 74.53 83.21
+CaFO-F*[30] 91.72 97.28 97.97 87.15 54.82 71.34 93.89 83.85 86.10 76.62 74.48 83.20
+Reprompt+Synthetic data 91.95 96.80 98.29 87.23 55.21 76.18 93.92 87.55 87.34 77.39 74.23 84.19
 Figure 6: Visualization of different learned visual feature manifolds via t-SNE.
 From left to right, we have CLIP, VLPT, and RePrompt.
-The image features extracted by RePrompt are more discriminating along the z-axis.
+The image features extracted by RePrompt discriminate more along the z-axis.
 Component ablation.
-We investigate the effectiveness of RePrompt and report results in Table 5.
+We investigate the effectiveness of RePrompt and report results in Table V.
 A progressive approach of incorporating retrieval modules is employed to demonstrate the efficacy of enhancement measures at various stages.
-The average accuracy of RePrompt consistently improves as we incrementally introduce retrieval augmentation: Retrieval-guiding training denoted as $+$Rg training loss, Retrieval-enhanced visual prompt denoted as $+$Re visual prompt, and Retrieval-based adapter denoted as $+$Rb adapter.
+The average accuracy of RePrompt consistently improves as we incrementally introduce retrieval augmentation: Retrieval-guided training denoted as $+$Rg training loss, Retrieval-enhanced visual prompt denoted as $+$Re visual prompt, and Retrieval-based adapter denoted as $+$Rb adapter.
 We also evaluated the Tip-adapter, which solely utilizes the retrieved samples as prototypes.
-The findings indicate that retrieval-enhanced visual prompts significantly enhance retrieval-based classification results.
+The findings indicate that retrieval-enhanced visual prompts significantly enhance retrieval-based classification.
 […]
@@
 The limited ability of Tip-Adapter-F to manipulate query features significantly hampers its performance.
-To further substantiate these observations, we employ t-SNE tSNE to visualize the 3D manifold of test features extracted from the CLIP, VLPT, and RePrompt models on the EuroSAT dataset.
-The t-SNE visualization results, presented in Figure 7, clearly depict that in the high-dimensional classification space, RePrompt achieves a much more distinct separation of image features, with feature distributions more aligned along the z-axis, distinguishing between various categories.
+To further substantiate these observations, we employ t-SNE [73] to visualize the 3D manifold of test features extracted from the CLIP, VLPT, and RePrompt models on the EuroSAT dataset.
+The t-SNE visualization results, presented in Fig.
+6, clearly depict that in the high-dimensional classification space, RePrompt achieves a much more distinct separation of image features, with feature distributions more aligned along the z-axis, distinguishing between various categories.
 Conversely, Tip-Adapter-F, constrained by a frozen visual encoder, exhibits t-SNE visualization results that closely resemble those of CLIP, with considerable overlap in feature distribution.
@@
 To this end, we conduct a supplementary experiment on ImageNet, where we substitute the convolution layer in ReConv with different layers.
-ReRNN is a variant in which the convolution layer is replaced with an LSTM layer LSTM ; RNN .
+ReRNN is a variant in which the convolution layer is replaced with an LSTM layer [74, 75].
 ReMLP replaces the convolution layer with a multi-layer perceptron layer.
 Additionally, the quality of high-retrieval samples also contributes to performance; hence, we examine the effect of utilizing original features and random sample features as inputs.
-ReConv($z_{q}$) is a variant that exclusively uses $z_{q}$ as input, whereas ReConv+Random employs random sample features.
+ReConv($z_{q}$) is a variant that exclusively uses $z_{q}$ as input.
+The ReConv+Random employs random sample features from the retrieval database.
 The experimental setup for these variants aligns with that of ReConv, maintaining a comparable number of parameters across both models.
-As demonstrated in Table 6, ReConv marginally outperforms the other variants across all shot levels.
+As demonstrated in Table VI, ReConv marginally outperforms the other variants across all shot levels.
 These results suggest that the specific network structure used for token and query fusion may not significantly affect the overall model performance.
@@
 However, the quality of retrieval samples plays a crucial role, implying that expanding the search library and enhancing the search methodologies are of greater importance.
+TABLE VI: Ablation study of the visual prompt learner across different input strategies and network structures.
+The first two rows showcase results using RNN and MLP structures, respectively.
+The middle two rows illustrate the impact of using original image features and random sample features as inputs.
+ReConv consistently outperforms the other variations, demonstrating its effectiveness across varying shot scenarios.
 method 16-shot 8-shot 4-shot 2-shot 1-shot
@@
 ReConv 74.53 73.36 71.84 70.68 69.87
-Table 6: Ablation study of the visual prompt learner across different input strategies and network structures.
-The first two rows showcase results using RNN and MLP structures, respectively.
-The middle two rows illustrate the impact of using original image features and random sample features as inputs.
-ReConv consistently outperforms the other variations, demonstrating its effectiveness across varying shot scenarios.
+TABLE VII: Comparison of training efficiency for different methods on 16-shot ImageNet. All experiments are trained with batch 16 on one RTX309 GPU.
 […]
-Overall, these results highlight the trade-offs between training duration, model complexity, computational efficiency, and performance, emphasizing the effectiveness of RePrompt in leveraging retrieval for improved few-shot classification on high-complexity tasks.
-[…]
-CoOP CoOP 71.20 0.4 14h40min 200
-Tip-Adapter-F tip-adapter 73.32 8.19 5min 20
+CoOP[6] 71.20 0.4 14h40min 200
+Tip-Adapter-F[21] 73.32 8.19 5min 20
 VLPT 71.42 0.45 15h50min 100
@@
 RePrompt 74.53 11.11 4h 20
-Table 7: Comparison of training efficiency for different methods on 16-shot ImageNet. All experiments are trained with batch 16 on one RTX309 GPU.
+TABLE VIII: Comparison of inference efficiency for different methods on 16-shot ImageNet. All experiments are trained with batch 16 and tested with batch size 32 on one RTX309 GPU.
 […]
-CoOP CoOP 71.20 299.64 162.5
-Tip-Adapter-F tip-adapter 73.32 10.5 42.5
+CoOP*[6] 71.20 299.64 162.5
+Tip-Adapter-F*[21] 73.32 10.5 42.5
 VLPT 71.42 214.84 162.5
@@
 RePrompt 74.53 57.60 76.5
-Table 8: Comparison of inference efficiency for different methods on 16-shot ImageNet. All experiments are trained with batch 16 and tested with batch size 32 on one RTX309 GPU.
-6 Retrieval Discussions
+[…]
+Overall, these results highlight the trade-offs between training duration, model complexity, computational efficiency, and performance, emphasizing the effectiveness of RePrompt in leveraging retrieval for improved few-shot classification on high-complexity tasks.
+VI Retrieval Discussions
 Given that the retrieval database is constructed from training data in a few-shot setting, the performance enhancements observed with RePrompt are intimately linked to data statistics.
 To delve deeper, we investigate the correlation between the efficacy of retrieval enhancement and the intra-class variance of visual features.
-This specific relationship is depicted in Figure 8, illustrating the association between intra-class visual variance and performance improvement.
+This specific relationship is depicted in Fig.
+7, illustrating the association between intra-class visual variance and performance improvement.
 Aside from the outliers—“StanfordCars” and “DTD”—the remaining datasets exhibit a discernible pattern.
@@
 This suggests a sensitivity of the RePrompt’s efficacy to the homogeneity of visual features within classes.
-6.1 Analysis of retrieval
-Experiments are conducted across FGVC Aircraft, Oxford Pets and ImageNet-1K to quantitatively assess the influence of retrieval in various few-shot scenarios.
-FGVC Aircraft has the smallest intra-class visual variance, while Oxford Pets has the largest; ImageNet-1K falls intermediate between these two.
-By comparing the optimal retrieval parameter adjustments among these datasets, we can analysis the influence of intra-class visual variance on the retrieval mechanism.
-The increasing value of the retrieval-enhanced prompt number $k_{\mathrm{re}}+2$ indicates that RePrompt benefits from additional retrieved information.
-As illustrated in Table 9, the optimal $k_{\mathrm{re}}$ parameter varies among the three datasets.
-Specifically, we observe that $k_{\mathrm{re}}=14$ consistently underperforms across all few-shot settings for Oxford Pets.
-This suggests that the retrieved information is detrimental and lacks confidence in contexts with significant intra-class visual variance.
-In contrast, $k_{\mathrm{re}}=14$ demonstrates improved average performance on FGVC Aircraft, underscoring its utility in scenarios with lesser intra-class variance.
+TABLE IX: Ablation study on different visual prompt numbers over Fgvc Aircraft and Oxford Pets with few-shot settings.
+The retrieval-enhanced prompt is mainly affected by the convolution kernel size of REConv.
 $k_{\mathrm{re}}$+2 Dataset 16-shot 8-shot 4-shot 2-shot 1-shot Ave
@@
 14+2 74.40 73.11 71.29 70.35 69.64 71.76
-Table 9: Ablation study on different visual prompt numbers over Fgvc Aircraft and Oxford Pets with few-shot settings.
-The retrieval-enhanced prompt is mainly affected by the convolution kernel size of REConv.
-The factor $n$ determines the number of retrieval samples, $\left|\boldsymbol{K}\right|=C\times\boldsymbol{n}$, used in retrieval-guiding training loss.
-Increasing the value of $n$ signifies a more relaxed constraint on retrieval.
-As shown in Table 10, $\boldsymbol{n}=1$ achieves optimal performance in the 16-shot setting on Oxford Pets, which implies that RePrompt prefers the output of CLIP with retrieval-enhanced prompt.
-However, in low-shot scenarios(e.g., $\boldsymbol{n}=2/1$ under 2/1-shot setting ), RePrompt may require more additional references due to the scarcity of training data.
+TABLE X: Ablation studies on retrieving $n$ samples for each class over Fgvc Aircraft and Oxford Pets with few-shot settings.
+The retrieval-guided training is weak in low-shot settings, since the model may require reference on the $k$-nn classifier.
 $\boldsymbol{n}$ Dataset 16-shot 8-shot 4-shot 2-shot 1-shot
@@
 1 74.37 73.23 71.24 70.38 70.02
-Table 10: Ablation studies on retrieving $n$ samples for each class over Fgvc Aircraft and Oxford Pets with few-shot settings.
-The $k$NN-guiding training is weak in low-shot settings, since the model may require reference on the $k$-nn classifier.
-Figure 9: Tuning retrieval parameter results over 3 datasets under the 16-shot learning setting.
-From top to bottom, images report ablation experiments about tuning retrieval parameters on FGVC Aircraft, ImageNet, and Oxford Pets, respectively.
+VI-A Analysis of Retrieval
+Experiments are conducted across FGVC Aircraft, Oxford Pets and ImageNet-1K to quantitatively assess the influence of retrieval in various few-shot scenarios.
+FGVC Aircraft has the smallest intra-class visual variance, while Oxford Pets has the largest; ImageNet-1K falls intermediate between these two.
+By comparing the optimal retrieval parameter adjustments among these datasets, we can analyze the influence of intra-class visual variance on the retrieval mechanism.
+The increasing value of the retrieval-enhanced prompt number $k_{\mathrm{re}}+2$ indicates that RePrompt benefits from additional retrieved information.
+As illustrated in Table IX, the optimal $k_{\mathrm{re}}$ parameter varies among the three datasets.
+Specifically, we observe that $k_{\mathrm{re}}=14$ consistently underperforms across all few-shot settings for Oxford Pets.
+In contrast, $k_{\mathrm{re}}=14$ demonstrates improved average performance on FGVC Aircraft, underscoring its utility in scenarios with lesser intra-class variance.
+This suggests that the retrieved information is detrimental and lacks confidence in contexts with significant intra-class visual variance.
+The factor $n$ determines the number of retrieval samples, $\left|\boldsymbol{K}\right|=C\times\boldsymbol{n}$, used in retrieval-guided training loss.
+Increasing the value of $n$ signifies a more relaxed constraint on retrieval.
+As shown in Table X, $\boldsymbol{n}=1$ achieves optimal performance in the 16-shot setting on Oxford Pets, which implies that RePrompt prefers the output of CLIP with retrieval-enhanced prompt.
+In low-shot scenarios(e.g., $\boldsymbol{n}=2/1$ under 2/1-shot setting ), RePrompt may require more additional references due to the scarcity of training data.
 Based on the above analysis, we summarize two empirical rules as follows: 1) For a dataset with low or high intra-class visual variance, it is imperative to employ a retrieval model with correspondingly strong or weak capabilities to guarantee stable performance improvements across scenarios.
+For example, if the dataset has a large intra-class visual variation, we need to control the impact of the research branch on the results by limiting the number of retrieved samples and retrieval-enhanced prompt.
 2) In low-shot settings, the inference outcomes are dominantly influenced by the retrieval results.
-6.2 Tuning retrieval parameters
-The final prediction distribution in RePrompt is influenced by the hyperparameters $\gamma$ and $\lambda$, and the prompt depth $J$.
-Ablation experiments are conducted to investigate the impact of these hyperparameters on the final outcomes across different datasets.
-As illustrated in Figure 9, from top to bottom, ablation experiments are carried out on FGVC Aircraft, ImageNet, and Oxford Pets under the 16-shot learning setting.
-The experiments adhere to the default configurations, except for modifying the depth $J$, $\gamma$, and $\lambda$ parameters.
-The experiment data are averaged over the results of experiments on seeds 1,2,3.
-The reference configuration is marked by a gray vertical dotted line.
-[…]
-[…]
-$\gamma$ varies.
-The hyperparameters $\gamma$ in the guiding loss $\mathcal{L}$ controls the intensity of the retrieval enhancement.
-As evidenced by the right part of Figure 9, the result continues to improve as $\gamma$ decreases.
-These results demonstrate that lower values of $\gamma$ enhance the reliance on retrieval-enhanced features, leading to significant improvements in classification accuracy.
-However, for FGVC Aircraft, the highest accuracy is observed at $\gamma=10^{-1}$, which is caused by its smallest intra-class visual variance.
-In summary, the hyper-parameters associated with retrieval critically influence the efficacy of retrieval enhancement.
-These findings highlight the necessity of striking an optimal balance between incorporating valuable retrieved information and mitigating the potential introduction of noise during the learning process.
-We provide valuable insights that are instrumental for the design of related methods and establish a guiding framework for future research in this domain.
-7 Conclusion
-In this paper, we investigate retrieval-enhanced visual prompt learning for visual-language models.
-Specifically, we propose integrating the results from a retrieval cache model into model inference using two approaches: prompts and predicted distributions.
+VII Conclusion
+In this paper, we propose RePrompt, a retrieval-based framework that enhances the performance of visual-language modes on few-shot classification tasks.
+Specifically, we integrate the retrieved results from a retrieval cache model into model inference using two approaches: prompts and predicted distributions.
 Additionally, We introduce a prompt learner that dynamically adapts prompts based on retrieval results and an auxiliary classification adapter.
@@
 Our study lays the groundwork for future research in enhancing visual-language models through retrieval mechanisms, and we encourage further investigations into these critical areas.
-8 Data Availability Statements
-The datasets that support the findings of this study are openly available in
-1) Few-shot image classification: ImageNet, Caltech101, OxfordPets, StanfordCars, Flowers102, Food101, FGVCAircraft, SUN397, DTD, EuroSAT, UCF101. Download split files from this link
-2) Domain generalization task: ImageNetV2, ImageNet-Sketch, ImageNet-A, ImageNet-R.
-3) Few-shot video understanding: HMDB51, UCF101, Something Something v2 (SSv2), Kinetics
-4) Few-shot Multi-few image classification: MVImgNet
-Acknowledgments
-This work was supported by National Key R&D Program of China (No. 2022ZD0118700) and Natural Science Foundation of Zhejiang Province (NO. LY21F030018).